### v1

# PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis

Abstract
We introduce PhysGaia, a novel physics-aware dataset specifically designed for Dynamic Novel View Synthesis (DyNVS), encompassing both structured objects and unstructured physical phenomena. Unlike existing datasets that primarily focus on photorealistic reconstruction, PhysGaia is created to actively support physics-aware dynamic scene modeling. Our dataset provides complex dynamic scenarios with rich interactions among multiple objects, where they realistically collide with each other and exchange forces. Furthermore, it contains a diverse range of physical materials, such as liquid, gas, viscoelastic substance, and textile, which moves beyond the rigid bodies prevalent in existing datasets. All scenes in PhysGaia are faithfully generated to strictly adhere to physical laws, leveraging carefully selected material-specific physics solvers. To enable quantitative evaluation of physical modeling, our dataset provides essential ground-truth information, including 3D particle trajectories and physics parameters, e.g., viscosity. To facilitate research adoption, we also provide essential integration pipelines for using state-of-the-art DyNVS models with our dataset and report their results. By addressing the critical lack of datasets for physics-aware modeling, PhysGaia will significantly advance research in dynamic view synthesis, physics-based scene understanding, and deep learning models integrated with physical simulation—ultimately enabling more faithful reconstruction and interpretation of complex dynamic scenes. Our datasets and codes are available in the project website [^1] http://cvlab.snu.ac.kr/research/PhysGaia.
1 Introduction
(a) Liquid with FLIP solver: hanok (left) and ship (right)
(b) Gas with Pyro solver: box-smoke (left) and pisa (right)
(c) Viscoelastic substances with MPM solver: pancake (left) and jelly party (right)
(d) Textile with Vellum solver: lucy (left) and basin (right)
Figure 1: Examples from the proposed physics-aware dataset, PhysGaia. They exhibit complex physical interactions between multiple objects composed of diverse materials such as liquid, gas, viscoelastic substance, and textile. This dataset will foster physics reasoning in dynamic scenes.
Since the introduction of Neural Radiance Fields (NeRF) [1], the Novel View Synthesis (NVS) algorithms based on deep learning have progressed rapidly. While early research primarily targeted static scenes, recent efforts have extended the scope to dynamic scene reconstruction to better support diverse and interactive AR/VR applications. This emerging direction, known as Dynamic Novel View Synthesis (DyNVS), aims to reconstruct 4D scenes from input videos and synthesize photorealistic images at novel viewpoints and time steps that are not seen during training.
The evolution of DyNVS has been closely tied to the development of suitable datasets. Early DyNVS datasets [2, 3, 4] primarily provided multiview training videos and typically involved only limited object motion. These datasets laid the foundation for early DyNVS research [3, 5, 6, 7, 8, 9, 10, 11], which focused on modeling scene dynamics by deforming canonical geometries over time. Recently, to better support AR/VR applications, datasets captured using handheld mobile devices have been introduced [7, 12, 13]. These datasets consist of monocular videos, and recent research has primarily focused on mitigating overfitting under such limited supervision. Building on these recent datasets, subsequent works [14, 15, 16, 17, 18, 19, 20] also have commonly extended recent Gaussian Splatting [21] to dynamic scenarios, aiming to achieve photorealistic rendering. While these efforts have led to significant improvements in visual quality, photorealism alone is insufficient for presenting the underlying dynamics and physical properties of real-world scenes.
As DyNVS continues to advance, a natural next step is to move beyond photorealism and incorporate physical realism—enabling models not only to render how scenes look, but also to reason about how they behave. Recent pioneering works [22, 23, 24, 25, 26] have begun to explore this direction by integrating differentiable physics simulation into Gaussian Splatting [21] frameworks. Despite this growing interest, current research largely remains limited to 4D generation [27, 28, 29, 30] or simplified DyNVS scenarios [26, 31], often restricted to single object or single material. Consequently, more complex scenarios—such as multi-object interactions or the modeling of diverse physical materials—are still significantly under-explored. In response, we introduce PhysGaia, a new benchmark designed to support and accelerate research in these emerging and challenging directions.
Our contributions are summarized as follows:
  • We introduce PhysGaia, a physics-aware dataset featuring rich interactions among multiple objects and encompassing a wide range of physical materials, including liquids, gases, textiles, and viscoelastic substances, as illustrated in Figure 1.
  • PhysGaia provides essential ground-truth information, such as 3D particle trajectories and physical parameters, enabling quantitative evaluation of physical modeling.
  • We evaluate existing state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and demonstrating the potential for improvement in this field.
The rest of this paper is organized as follows. Section 2 reviews related work and highlights the uniqueness of our dataset. We discuss the main properties and potential research enabled by our dataset in Section 3. Section 4 details material-specific physics solvers used for dataset construction and methodologies for generating multi-body interaction scenarios. Section 5 shows our analysis, including evaluations of recent DyNVS algorithms on PhysGaia, We conclude the paper in Section 6.
2 Related Work
2.1 Dynamic Novel View Synthesis
In recent years, significant advances have been made in novel view synthesis [1, 32, 33, 34, 35, 21]. Although this field initially have focused on static scene reconstruction but now it extended to handling dynamic scenes, which is known as Dynamic Novel View Synthesis (DyNVS). Early DyNVS methods were built on Neural Radiance Fields (NeRF) [5, 6, 7, 8, 9, 10, 11], which usually modeled scene dynamics either by implictly modeling with temporal inputs [5, 6] or directly estimating the time-wise deformation of canonical geometry through auxiliary neural networks [7, 8, 9, 10, 11]. Following the emergence of 3D Gaussian Splatting (3DGS) [21], recent DyNVS research [17, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] has shifted toward Gaussian Splatting-based representations to achieve higher fidelity and real-time inference. Since Gaussian Splatting is an explicit representation, additional deformation networks can directly modify the scale, rotation, and position of each gaussian primitive over time.
In a significant development, research is now moving towards constraining the deformation of gaussian primitives to adhere to physical realism. PhysGaussian [22] pioneered this by combining an MPM (Material Point Method) [48] simulator with Gaussian Splatting [21], treating each gaussian primitive as a particle within the MPM simulation. This work has inspired many subsequent studies [31, 26, 27, 28, 29, 30] integrating physics-aware priors into Gaussian Splatting; however, these efforts remain largely confined to generation tasks [27, 28, 29, 30], with only a few addressing Dynamic Novel View Synthesis (DyNVS) [31, 26]. Even for DyNVS, methods are typically limited to single objects or single materials, with most focusing on viscoelastic substances. Consequently, exploring physics-aware DyNVS research involving rich object interactions and diverse physical materials remains significantly underexplored. We believe our dataset can serve as a foundational resource for this crucial direction.
2.2 4D Datasets for Dynamic Novel View Synthesis
Table 1: Comparison with Dynamic Novel View Synthesis datasets. This table summarizes key attributes including the number of total scenes having test videos for DyNVS evaluation, the number of scenes with physical phenomena, presence of multi-body interaction, ground-truth physics parameters, e.g., viscosity, and ground-truth 3D particle trajectories. Our PhysGaia dataset uniquely offers a comprehensive set of diverse scenes with complex multi-body interactions and complete ground-truth physical information, a significant advancement over prior benchmarks.
Datasets Scene Stats Physics Information Capture Type
# total scene # physics Interaction Physics param. 3D Traj.
Plenoptic [2]   6       1(gas) No No No Multiview
D-NeRF [3]   8       1(viscoelastic) No No No Monocular
NVIDIA Dynamic [4]   8       0 No No No Monocular
Nerfies [7]   4       0 No No No Monocular
HyperNeRF [12]   4       0 No No No Monocular
DyCheck [13]   7       1(textile) No No No Monocular
NeRF-DS [49]   8       0 No No No Monocular
EvDNeRF [50]   6       0 No No No Multiview
Synthetic Soccer [51]   3       0 No No No Multiview
HDR-HexPlane [52]   8       0 No No No Multiview
[…]
Table 1 provides a comparative overview of our dataset alongside existing multiview and monocular DyNVS datasets. The initial DyNVS datasets [2, 3, 4] primarily employed multiview configurations or captured scenes with very limited motion, typically involving mostly rigid objects. For example, Plenoptic [2] was captured with sufficiently many multiview cameras at fixed positions. To advance DyNVS toward practical AR/VR applications, more user-friendly datasets, often captured using handheld mobile phones, were later introduced. Nerfies [7] pioneered handheld iPhone captures, though its scenes remained largely static. HyperNeRF [12] then introduced more rapid and varied motions, and DyCheck [13] further resolved camera teleportation issues seen in HyperNeRF. These monocular datasets have motivated research mitigating overfitting to training videos in these settings. Some approaches [14, 15, 16] focus on leveraging additional priors such as diffusion models [53], depth estimation models [54], and point trackers [55], while others [15, 17, 18, 19, 20] emphasize constraining deformation by adopting techniques like motion factorization [56] or As-Rigid-As-Possible (ARAP) regularization [57]. Nonetheless, the primary objective across these datasets remains photorealistic reconstruction, with limited emphasis on physics-aware dynamic modeling  [^2] Note that some recent datasets have also been introduced, but they are typically tailored to specific scenarios, such as event cameras [50], HDR rendering [52], specular lighting effects [49], or human motion capture [51]..
Unlike existing DyNVS datasets, our PhysGaia dataset offers scenes with multi-body interactions involving diverse physical materials, as detailed in Table 1. While some datasets [2, 3, 13] include limited scenes with physical phenomena, these are typically few in number, not generated using accurate physics solvers, and lack complex multi-object interactions. In this context, PhysGaia occupies a unique position and holds strong potential to spur advancements in physics-aware dynamic scene modeling and reasoning.
2.3 4D Datasets from Physics Simulator
Table 2: Comparison with physics-simulated datasets. “Multi-body Inter.” denotes multi-body interaction. Unlike existing datasets, our PhysGaia offers rich interactions, a diverse range of physical materials, and access to accurate physics parameters and ground-truth 3D particle trajectories, making it a uniquely valuable resource for advancing physics-based understanding of dynamic scenes.
Datasets Multi-body Inter. Materials Physics Information
Liquids Gas Textile Viscoelastic Physics params. 3D traj.
DNG [58] Yes $\checkmark$ No Yes
CLOTH4D [59] Yes $\checkmark$ No Yes
4D-DRESS [60] Yes $\checkmark$ No Yes
Rasheed et al. [61] No $\checkmark$ Yes No
Deng et al. [62] No $\checkmark$ No No
PAC-NeRF [63] No $\checkmark$ $\checkmark$ Yes Yes
Spring-Gaus [31] No $\checkmark$ Yes Yes
ScalarFlow [64] No $\checkmark$ Yes Yes
[…]
(a) PAC-NeRF [63]
(b) Spring-Gaus [31]
(c) ScalarFlow [64]
Figure 2: Visualization of datasets most similar to our PhysGaia. While all of these datasets address physical phenomena, they are limited in several key aspects: limited coverage of physical materials, overly simplified dynamics, and an absence of rich multi-body interactions.
To provide a comprehensive comparison, we evaluate our PhysGaia against other 4D datasets generated via physics simulation, as summarized in Table 2. Since DyNVS tasks require multiview RGB imagery, we focus on datasets that offer such data. While some existing datasets support multi-body interactions, they often lack rich interaction dynamics or are confined to a single material type. For instance, clothed human datasets [58, 59, 60] exhibit textile-human interactions but are limited to textiles and human-centric motions. In terms of ground-truth physical information, a few datasets [63, 31, 64] provide both physical parameters and 3D trajectories. However, these datasets are typically restricted to single materials and do not capture complex, multi-body interactions, as illustrated in Figure 2. Specifically, PAC-NeRF [63] (Figure 2(a)) focuses on liquids and viscoelastic materials, but its liquid scenarios are highly viscous, behaving similarly to viscoelastic materials. Spring-Gaus [31] (Figure 2(b)) is restricted to viscoelasticity, and ScalarFlow [64] (Figure 2(c)) is limited to non-interacting gases.
In stark contrast, our PhysGaia directly addresses these limitations. It provides rich interactions, a diverse range of physical materials, and access to accurate simulation parameters and ground-truth trajectories. This makes PhysGaia a uniquely valuable dataset for advancing our understanding of physics in dynamic scenes.
3 Dataset Properties and Research Impact
We propose PhysGaia to advance physically realistic reconstruction in Dynamic Novel View Synthesis (DyNVS), extending the conventional reconstruction task beyond photorealism. As highlighted in Tables 1 and 2 of Section 2, our dataset uniquely features complex multi-body interactions involving a diverse range of physical materials, distinguishing it within the DyNVS landscape. We believe this comprehensive dataset holds significant potential to enhance the understanding of physics in dynamic scenes. This section outlines the key properties of our dataset and discusses potential future research directions it enables.
3.1 Dataset Properties
Complex physics-aware dynamics
(a) Multi-body interaction: Jelly party
(b) Multi-body interaction: Ship
(c) Splashing (non-locally rigid)
(d) Specular modeling
(e) Refraction
Figure 3: Visualization of physics properties in PhysGaia. Alongside multi-body interactions, PhysGaia also includes various physical phenomena like splashing, refraction, and specular effects.
Our PhysGaia dataset comprises 17 scenes that include multi-body interactions. Figure 1 illustrates a subset of these scenes. We further visualize the detailed examples in Figure 3(a) and Figure 3(b), which exhibit collisions between viscoelastic objects causing rapid changes in motion direction, and rigid-liquid surface impacts resulting in splashing, respectively. To ensure that these scenes adhere to physical laws with accurately calculated force exchange among multi bodies, we carefully select material-specific solvers: FLIP for liquids, Pyro for gases, MPM for viscoelastic materials, and Vellum for textiles. Further details on simulation configurations for handling multi-body interactions are found in Section 4.
Beyond multi-body interactions, our dataset also exhibits various physical phenomena such as non-locally rigid motion (common in liquid and gas scenes), specular reflection, and refraction, as shown in Figures 3(c)3(d), and 3(e), respectively. These properties enhance the realism of our dataset and enable a wide range of downstream tasks and research applications.
Ground-truth physics information
In contrast to real-world video datasets, which are easier to capture but lack access to underlying physical properties, our simulated dataset provides complete ground-truth physical information. This includes 3D particle trajectories and accurate physics parameters such as viscosity, Young’s modulus, Poisson’s ratio, and temperature for gas-based scenes [^3] Temperature is a crucial component for buoyancy calculation in smoke simulations.. This comprehensive provision enables precise evaluation of physical reasoning in dynamic scenes and directly supports the future research directions discussed in Section 3.2.
Supporting diverse DyNVS tasks
Our dataset uniquely supports both multiview and monocular DyNVS task. Unlike most multiview datasets [2, 50, 51, 52] with videos captured from fixed camera positions like CCTV, PhysGaia provides several moving monocular video sequences with independent trajectories as training data. This allows for diverse training configurations; using full sequences corresponds to a multiview DyNVS task, while using a single sequence corresponds to a monocular DyNVS task. The monocular setup matches datasets like DyCheck [13], providing realistic handheld camera inputs without synthetic teleportation. For evaluation, we employ two static cameras with a large baseline.
Customizability
(a) RGB
(b) Depth
(c) Normal
(d) Re-lighted
Figure 4: Examples of diverse modalities that users can generate from the provided simulation node graphs.
We provide the complete simulation node graphs and their exact parameter settings used to construct our dataset. These graphs encompass all relevant components, including physics solvers, source geometries, camera positions, lights, materials, and texture controls. By modifying these nodes, users can easily generate customized scenes or additional modalities such as depth maps, surface normals, and re-lighted images tailored to their specific downstream tasks, as shown in Figure 4.
Accessibility
Our PhysGaia benchmark is designed for research-friendly and easy access, providing integration pipelines that enable the use of state-of-the-art DyNVS models [15, 36, 37, 38] with our data, as detailed on our GitHub repository [^4] https://github.com/mjmjeong/PhysGaia. Furthermore, we include COLMAP-reconstructed point clouds for each scene. These provisions aim to foster the adoption of our dataset by researchers working for state-of-the-art DyNVS models.
3.2 Potential Research
We highlight the potential impact of our PhysGaia dataset by outlining several promising future research directions it uniquely enables.
Advancing physical reasoning in dynamic scenes
Since our PhysGaia dataset provides ground-truth physics information, it facilitates precise evaluation of physical reasoning in dynamic scenes. For instance, ground-truth physics parameters like viscosity can be used to evaluate inverse physics estimation methods, where differentiable simulators are employed to optimize these parameters. Furthermore, unlike existing DyNVS research that primarily focuses on photorealism and relies solely on ground-truth RGB images, our dataset offers ground-truth 3D trajectories. This enables evaluating the actual motion of individual Gaussian primitives in recent GS-based methods, moving beyond just visual fidelity. We believe this unique feature establishes our dataset as a valuable benchmark for developing and evaluating physics-aware DyNVS models./asset
Exploring multi-body interactions
While recent research integrates physics into DyNVS [26, 31], it largely remains limited to single materials and often single objects within scenes. As a result, the crucial aspect of physics reasoning for interactions between multiple objects in DyNVS—particularly for the estimation of force exchange and deformation during contact—remains largely unexplored. We believe our novel dataset—specifically designed to capture complex multi-body interactions—will serve as a valuable resource for advancing future research in multi-physics modeling and adaptive representations for hybrid scenes.
Integrating material-specific physics solvers
The dominant approach to incorporating physics into DyNVS algorithms currently relies on differentiable particle-based simulators [48, 65, 66, 67, 68, 69], which treat Gaussian primitives as particles. However, different physical phenomena are best captured by different physics solvers—a principle reflected in our dataset construction process described in Section 4. This insight can guide researchers seeking to integrate more appropriate solvers tailored to specific material behaviors, such as fluids, cloth, and smoke. For instance, the FLIP [70] solver excels in simulating incompressible fluids due to its hybrid particle-grid representation, offering improved stability and realism compared to purely particle-based SPH-based solvers. Similarly, thermodynamic phenomena, such as smoke dynamics, are conventionally handled by voxel-based simulators. Nevertheless, the seamless integration of these volumetric solvers into particle-based representations like Gaussian Splatting presents a significant and largely unexplored research challenge.
4 Dataset Construction
Our dataset consists of 17 scenes, categorized into four material types: liquids, gases, viscoelastic materials, and textiles. Each category contains 4 to 5 scenes. This section details the solver properties for each material and provides key simulation details regarding multi-body interactions. Please refer to our supplemental document for other simulation configurations.
4.1 Common Setup
Simulator selection
We select SideFX Houdini 20.5 as the foundation of our physics-informed data-generation pipeline because it integrates multiple physics solvers within a unified procedural environment. By sharing a common computational graph, it ensures consistent multi-body interactions under uniform boundary conditions. We access simulation data, such as particle positions and flow fields, on a per-frame basis via its Python API.
Rendering
All frames are rendered at a resolution of $640\times 720$. We use path tracing with 256 samples per pixel and apply NVIDIA OptiX for denoising. Scenes are illuminated with 1–3 point lights (intensity: 600–4000); shadows are disabled in textile-focused scenes like tube-flag to emphasize geometry. All cameras follow spiral trajectories to capture diverse views over time.
4.2 Liquid
For liquid scenes, we adopt the Fluid-Implicit Particle (FLIP) solver [70], a hybrid particle-grid method. FLIP maintains particle velocities throughout the simulation, using the grid solely to compute and apply forces such as pressure and viscosity. This approach preserves fine-scale, high-frequency particle velocities, which is crucial for modeling realistic and rapid liquid behavior while adhering to the Navier–Stokes equations. While the Material Point Method (MPM) [48] can also model fluids, its direct velocity aggregation onto the grid limits its ability to capture highly dynamic fluid phenomena like splashing. Particle-only solvers [71] are another option, but FLIP is generally better suited for incompressible fluids due thanks to its hybrid grid-based representation.
Dynamic interactions
For scenes where fluid spills onto fixed objects, such as ice and hanok, we employ a surface operator to simulate multi-body interaction. This approach is computationally efficient as interaction primarily occurs near the surface. In contrast, for scenes like ship and cereal where objects fall into liquid, causing both fluid and objects to move and influence each other through force exchange, we use a dynamic operator to accurately model these more complex interactions.
4.3 Gas
For gas (smoke) simulation, we utilize the Pyro solver [72]. Pyro accurately models the temperature field, which is essential for capturing buoyancy effects in gaseous materials. It employs grid-based representations of density, velocity, and temperature, ensuring compliance with the Navier–Stokes equations governing fluid mechanics. Since ground-truth motion is represented as a velocity field and storing full velocity fields can require up to 2GB per frame, we provide a subsampled set of particle trajectories per scene to facilitate efficient data storage and processing.
Dynamic interactions
In the pisa scene, we reduced the voxel size from the default 0.1 to 0.05 to better capture the tower’s intricate details and added a lateral wind of speed 2 to wrap the plume around it. Other scenes involving gaseous materials utilized the default simulation settings.
4.4 Viscoelastic Substances
We use MPM [48] for viscoelastic substances like snow and jelly. As an extension of FLIP [73], MPM is ideal for simulating chunk-based materials. It aggregates particle information onto a grid, performs computations, and then reprojects results back to particles, effectively capturing deformation and internal force propagation.
Dynamic interactions
For the pancake scene, we reduced the grid size from the default 0.025 to 0.002 to more faithfully capture its thin-sheet dynamics; other scenes use the default grid size. To suppress spurious artifacts that can arise from aggregating particle properties onto the grid, we increased the number of samples participating in node calculations by oversampling. The oversampling scales are set to 6, 2, 4, and 2 for the bouncing balls, cow, jelly party, and pancake scenes, respectively. Additionally, for the bouncing balls scene, we added a static bowl-shaped collider so that the falling balls rebound off both one another and the bowl’s surface.
4.5 Textile
For textile materials, we adopt the Vellum solver [74], which is based on the Extended Position Based Dynamics (XPBD) framework [75]. XPBD improves upon classical Position Based Dynamics (PBD) [76] by integrating a Lagrange multiplier and its update. This effectively decouples material stiffness from the solver’s time-step size and iteration count, making Vellum a widely used method for simulating deformable objects, especially cloth.
Dynamic interactions
To simulate interactions between objects and textiles–where both move and exchange forces, as seen particularly in the basin scene–we employ the shape match constraint. This constraint helps maintain the overall shape of objects by driving points toward their rest configuration, allowing the material to preserve its structural integrity while still interacting dynamically with textiles and other objects. For the lucy scene, we increased the simulation sub-step count fivefold over the default to robustly handle collisions with the statue’s complex geometry. To simulate wind-induced fluttering in the flags, single-flag, and tube scenes, we applied external forces using a POP Wind node that blows parallel to the ground plane.
5 Analysis
5.1 Evaluation of Existing Algorithms on PhysGaia
Experimental setup
We evaluate recent DyNVS methods on our PhysGaia dataset: D-3DGS [36], 4DGS [37], STG [38], and SOM [15]. Except for SOM [15], which is specialized for monocular setups, all models were tested under both monocular and multiview setups. For point cloud initialization, we first run COLMAP for dense matching and fusion, then uniformly downsample its output to approximately 40,000 points. All experiments are conducted on a NVIDIA A6000 GPU. For evaluation, we adopt standard image quality metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [77], and Learned Perceptual Image Patch Similarity (LPIPS) [78].
Quantitative results
Table 3: Average quantitative results for each material category across all baseline algorithms. Performance tends to be higher for textile materials, but it notably worsens for viscoelastic materials, which typically exhibit highly complex motions involving numerous independent objects.
Capture Type Method Liquid Gas
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Monocular D-3DGS [36] 22.7 0.87 0.22 21.9 0.89 0.16
4DGS [37] 24.2 0.87 0.23 21.7 0.88 0.17
STG [38] 19.2 0.72 0.39 21.9 0.85 0.24
SOM [15] 19.6 0.80 0.32 20.0 0.84 0.27
Multiview D-3DGS [36] 22.2 0.87 0.24 23.7 0.91 0.13
4DGS [37] 25.1 0.88 0.22 24.2 0.89 0.17
STG [38] 20.8 0.75 0.40 25.0 0.91 0.19
Capture Type Method Viscoelastic materials Textile
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Monocular D-3DGS [36] 20.1 0.84 0.15 22.1 0.83 0.18
4DGS [37] 19.5 0.82 0.18 24.9 0.84 0.18
STG [38] 13.6 0.63 0.40 21.9 0.84 0.21
SOM [15] 16.7 0.75 0.23 20.7 0.79 0.22
Multiview D-3DGS [36] 22.2 0.89 0.10 27.7 0.90 0.12
4DGS [37] 21.0 0.85 0.15 26.6 0.87 0.15
STG [38] 17.2 0.70 0.36 21.1 0.81 0.25
Tables 3 and 4 report average performance by material type and across all 17 scenes. As expected, multiview setups outperform monocular ones, yet even they often yield PSNR below 30. This underscores the difficulty of capturing complex multi-body dynamics in our dataset, suggesting physics-aware priors are essential beyond RGB-based fitting alone. Among materials, textiles are generally reconstructed more accurately, while viscoelastic materials pose greater difficulty due to their complex, object-rich motion. Interestingly, SOM [15], despite strong results on other datasets, performs inconsistently on ours. We presume its motion factorization struggles with highly dynamic phenomena such as splashing and multiple interacting objects. Additionally, its heavy reliance on external modules like point trackers leads to performance drops when these modules fail in challenging scenes, particularly those lacking distinct visual features like smoke or gas.
Qualitative results
(a) GT image
(b) D-3DGS [36]
(c) 4DGS [37]
(d) STG [38]
(e) SOM [15]
Figure 5: Qualitative results of recent DyNVS methods on the jelly party, lucy, and hanok scenes with monocular setup. All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
Figure 5 presents qualitative results from monocular settings, revealing that all methods struggle to accurately capture multi-body interactions. This leads to common issues like needle-like artifacts and under-reconstruction of dynamic regions. For 4DGS [37], its default grid learning rate often causes NaN loss values; while reducing it stabilizes training, the output resembles static scenes due to poor dynamic capture. More results are available in the supplementary material.
5.2 Limitation in Physical Realism
We analyze a key limitation of recent DyNVS methods regarding physical realism. Since Gaussian Splatting-based representation, used in recent algorithms, is an explicit representation, the motion of gaussian primitives is expected to closely follow the actual physical trajectories of dynamic scenes. However, all tested DyNVS algorithms consistently fail to accurately reconstruct these physically reliable motions, as shown in Figure 6. D-3DGS [36], 4DGS [37], and STG [38] produce only localized fluctuations, while SOM [15] exhibits erratic motion primarily due to inaccuracies from external point trackers. This discrepancy is acutely apparent in fluid or gas scenes. Here, photorealism is easily achieved via color similarity alone, allowing primitives to simply fluctuate locally near the surface rather than genuinely modeling true dynamics. This analysis highlights a critical limitation—relying solely on RGB reconstruction loss ensures photorealism but does not guarantee physical realism. We believe that our PhysGaia dataset, which provides ground-truth trajectories, will be instrumental in advancing future research toward physics-aware dynamic scene reconstruction.
Table 4: Average quantitative results across all 17 scenes for monocular and multiview settings. Multiview setups generally offer better reconstruction performance than monocular ones, yet PSNR scores remain below 30 even with multiple views. This highlights the inherent difficulty of reconstructing complex multi-body interactions within our dataset.
      Method       Monocular       Multiview
      PSNR $\uparrow$       SSIM $\uparrow$       LPIPS $\downarrow$       PSNR $\uparrow$       SSIM $\uparrow$       LPIPS $\downarrow$
      D-3DGS [36]       21.7       0.86       0.18       24.2       0.89       0.14
      4DGS [37]       22.7       0.85       0.19       24.4       0.87       0.17
      STG [38]       19.3       0.76       0.30       21.0       0.79       0.30
      SOM [15]       19.3       0.80       0.26       N/A       N/A       N/A
(a) GT image
(b) D-3DGS [36]
(c) 4DGS [37]
(d) STG [38]
(e) SOM [15]
Figure 6: Comparison of reconstructed trajectories and its ground truth on the Box-smoke and Pisa scenes. The reconstructed trajectories do not accurately follow the true physical trajectories.
6 Conclusion
We propose a novel physics-aware dataset, PhysGaia, specifically designed to understand physics in dynamic scenes, particularly for Dynamic Novel View Synthesis (DyNVS). Comprising 17 diverse scenes, our dataset captures complex multi-body interactions with a wide variety of materials. Each scene is faithfully generated using material-specific physics solvers, ensuring adherence to physical laws and providing rich ground-truth physics data including 3D particle trajectories and physics parameters. This ground truth data uniquely enables the evaluation of physics reasoning. We also test state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and highlighting significant potential for improvement. We believe PhysGaia will be a critical resource that accelerates progress in physics-aware dynamic scene understanding.
Appendix A Details of PhysGaia
A.1 Scene Lists
Our dataset consists of 17 scenes divided into four categories: liquid, gas, viscoelastic substances, and textile. Each category contains 4 to 5 scenes, and the detailed scene names are listed in Table 5.
Table 5: List of scenes included in our PhysGaia dataset
Liquid Gas Viscoelastic substance Textile
Scene name Cereal Pisa Jelly party Lucy
Ship Box-smoke Pancake Basin
Hanok Single smoke Bouncing balls Flags
Ice Falling Cow Single flag
Tube
A.2 Camera Trajectories
(a) $180^{\circ}$ scene: 2 cams
(b) $360^{\circ}$ scene: 4 cams
Figure 7: Visualization of training camera’s traj.
Figure 7 visualizes our training camera trajectories, where we adopt either 2 or 4 cameras depending on the properties of each scene. For the cow, single flag, and tube scenes–which resemble $180^{\circ}$ settings with background walls placed behind the objects–we use only 2 cameras, as shown in Figure 7(a). In contrast, for $360^{\circ}$ scenes, we place 4 cameras at evenly spaced viewpoints around the scene, enabling full coverage from four directions. All cameras follow spiral trajectories to capture diverse views over time. For the monocular setup, we select one of the training cameras to generate the training data.
A.3 Implementation Details
All simulations were conducted on 4 NVIDIA Titan Xp GPUs and Intel Xeon E5-2630 v4 CPUs with 40 cores. While most scenes rendered in 2–3 hours, the lucy and hanok scenes required 12–24 hours due to their complex geometry and increased simulation costs. To supplement Section 4, we provide additional detailed simulation implementation along with the value of physics parameters. These parameters can be served as ground-truth labels for downstream physical reasoning tasks.
Liquid
We assume incompressible, non-Newtonian fluids in our simulations. Table 6 summarizes the density and viscosity values, which are the most important factors governing fluid motion. External forces are limited to gravity. For the ship scene, we reduced the particle separation from 0.1 to 0.05 to decrease initial particle displacement and improve simulation accuracy. Also, the particle separation during dynamic simulation is set to 0.1 for the cereal scene and 0.05 for the ship scene.
Table 6: Physics parameters used for liquid materials.
      Scene name       Object       Density ($kg/m^{3}$)       Viscosity
      Ship       Water       800       0
      Ship       300       -
[…]
      Cereal       1000       -
[…]
      Ice       Water       1000       0
Gas
In gas simulations, temperature is the most critical physical parameter, governing buoyancy, expansion, and flow dynamics. We initialized the temperature to 3000K for gas materials. As temperature evolves locally due to gas movement and interactions, its spatiotemporal variation is represented as a temperature field defined on a voxel grid. These temperature fields are accessible via the provided simulation source file with Python API, similar to how flow fields are accessed for generating ground-truth trajectory data.
Viscoelastic substances
Gravity is applied to all scenes; however, for the cow scene, we introduce additional internal forces to induce cow motion and increase the complexity of the physical reasoning scenarios. Boundaries are set to be open in all directions, allowing objects to move freely without collisions against invisible walls. Table 7 lists detailed physics parameters, including Young’s modulus ($E$) and Poisson’s ratio ($P$), used in the simulations.
Table 7: Physics parameters used for viscoelastic materials.
Scene name Object Type $E$ $P$ Viscosity
Pancake Honey Viscous $2.5\times 10^{5}$ 0.10 0.125
Pancakes Chunky $8.0\times 10^{4}$ 0.23 -
[…]
Fishbowl Static - - -
[…]
Cow Cow Elastic $5.0\times 10^{4}$ 0.45 -
Textile
For textiles, we primarily adopt the default configuration from the Vellum solver, with detailed modifications summarized in Table 8.
Table 8: Physics parameters used for textile materials.
Scene name Object Stretch Bend
Stiffness Damping Ratio Stiffness Damping Ratio
Flags Flag1 (small) $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
Flag2 (big) $1.0\times 10^{11}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
[…]
Cloth $1.0\times 10^{13}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
[…]
Lucy Cloth $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
Tube Tube flag $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
A.4 Scene Visualization
Figure 8 and Figure 9 visualize more scenes in PhysGaia to supplement the Figure 1 in our main paper. As shown in Table 2 of the main paper, textile materials remain underexplored in physics-based datasets. To facilitate progressive research development in this domain, we include few simpler scenes such as single flag and tube that employ with only basic wind interactions. We believe this approach effectively supports advancements by providing more accessible starting points. Similarly, our dataset also includes few simpler gas scenes, simple smoke and falling, for the same reason as with textile scenarios: to facilitate progressive development. For the cow scene, internal forces intrinsic to the cow itself introduce unique and more complex dynamics, rather than using only gravity.
(a) Liquid: hanok (top-left), ship (top-right), cereal (bottom-left), ice (bottom-right)
Figure 8: Examples from the proposed physics-aware dataset, PhysGaia.
(a) Gas: box-smoke (top-left), pisa (top-right)), simple smoke (bottom-left), and falling (bottom-right)
(b) Viscoelastic substances: pancake (top-left), jelly party (top-right), bouncing balls (bottom-left), and cow (bottom-right)
(c) Textile: lucy (top-left), basin (top-right), flags (mid-left), tube (mid-right), and single flag (bottom)
Figure 9: Examples from the proposed physics-aware dataset, PhysGaia. They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, viscoelastic substance, and textile. This dataset will foster physics reasoning in dynamic scenes.
Appendix B Details of Experimental Results
To supplement Table 3, Table 4, and Figure 5 of the main paper, we report more detailed quantitative results along with additional qualitative results.
B.1 Performance breakdown: monocular setting
We provide per-scene breakdown performance of monocular video reconstruction setting in Table 9. We report the performance of D-3DGS [36], 4DGS [37], STG [38], and SOM [15], which serve as the most common and recent baselines for the DyNVS task.
Table 9: Per-scene breakdown results for all 17 scenes under the monocular setting.
Method Cereal Ship Hanok
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 23.3 0.88 0.15 25.9 0.90 0.16 16.3 0.78 0.28
4DGS [37] 26.3 0.90 0.14 25.9 0.91 0.15 15.5 0.74 0.32
STG [38] 15.8 0.54 0.54 22.8 0.87 0.23 14.8 0.60 0.41
SOM [15] 21.2 0.81 0.22 23.8 0.87 0.23 14.7 0.67 0.40
Method Ice Pisa Box-smoke
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 25.3 0.92 0.30 20.7 0.72 0.26 20.7 0.96 0.13
4DGS [37] 29.2 0.92 0.29 19.3 0.67 0.26 21.3 0.96 0.12
STG [38] 23.1 0.87 0.39 20.1 0.68 0.33 22.7 0.95 0.16
SOM [15] 18.7 0.84 0.41 17.1 0.65 0.41 23.8 0.95 0.18
Method Single smoke Falling Jelly party
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 26.5 0.97 0.08 19.6 0.90 0.18 16.3 0.81 0.21
4DGS [37] 26.1 0.97 0.08 20.1 0.90 0.19 14.9 0.70 0.28
STG [38] 24.5 0.97 0.10 20.2 0.77 0.37 11.1 0.51 0.44
SOM [15] 23.9 0.95 0.18 15.3 0.81 0.33 14.9 0.69 0.28
Method Pancake Bouncing balls Cow
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 22.8 0.88 0.13 19.8 0.77 0.15 21.6 0.92 0.10
4DGS [37] 17.0 0.78 0.22 22.7 0.86 0.10 23.6 0.92 0.11
STG [38] 10.6 0.60 0.39 11.9 0.51 0.60 20.6 0.87 0.16
SOM [15] 13.9 0.68 0.34 17.4 0.79 0.14 20.7 0.85 0.17
Method Lucy Basin Flags
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 22.8 0.92 0.10 18.2 0.67 0.36 23.6 0.94 0.12
4DGS [37] 27.5 0.94 0.08 18.0 0.68 0.38 31.9 0.96 0.08
STG [38] 19.6 0.86 0.19 16.8 0.66 0.43 24.7 0.90 0.17
SOM [15] 21.0 0.90 0.17 16.3 0.61 0.43 27.4 0.93 0.12
Method Single flag Tube Average
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 18.3 0.68 0.24 27.7 0.96 0.05 21.7 0.86 0.18
4DGS [37] 18.0 0.65 0.29 28.8 0.96 0.08 22.7 0.85 0.19
STG [38] 26.0 0.87 0.14 22.1 0.92 0.12 19.3 0.76 0.30
SOM [15] 16.3 0.60 0.24 22.5 0.93 0.15 19.3 0.80 0.26
B.2 Performance breakdown: multiview setting
We provide per-scene breakdown performance for the multiview video reconstruction setting in Table 10. We report results for D-3DGS [36], 4DGS [37], and STG [38], which serve as the most common and recent baselines for the DyNVS task. Note that since SOM [15] is specialized for monocular video setups, we omit its performance in the multiview evaluation.
Table 10: Per-scene breakdown results for all 17 scenes under the multiview setting.
Method Cereal Ship Hanok
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 28.1 0.93 0.11 28.8 0.93 0.11 15.6 0.76 0.31
4DGS [37] 27.6 0.92 0.13 25.6 0.91 0.15 15.8 0.75 0.33
STG [38] 15.8 0.58 0.59 25.9 0.90 0.20 15.7 0.65 0.43
Method Ice Pisa Box-smoke
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 16.4 0.86 0.43 22.3 0.75 0.23 20.5 0.96 0.11
4DGS [37] 31.4 0.93 0.28 22.8 0.74 0.24 22.6 0.97 0.07
STG [38] 25.9 0.89 0.38 28.7 0.90 0.18 19.0 0.95 0.17
Method Single-smoke Falling Jelly party
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 26.4 0.98 0.07 25.7 0.95 0.11 18.2 0.87 0.16
4DGS [37] 28.2 0.98 0.07 23.4 0.87 0.30 16.1 0.80 0.24
STG [38] 27.0 0.98 0.07 25.1 0.83 0.34 12.5 0.60 0.39
Method Pancake Bouncing balls Cow
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 25.3 0.90 0.08 25.7 0.91 0.07 19.8 0.87 0.10
4DGS [37] 18.7 0.84 0.18 25.0 0.84 0.09 24.1 0.92 0.10
STG [38] 11.7 0.65 0.36 12.9 0.59 0.54 31.8 0.95 0.15
Method Lucy Basin Flags
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 30.2 0.96 0.06 23.6 0.82 0.24 30.5 0.96 0.09
4DGS [37] 28.6 0.95 0.07 20.5 0.75 0.33 32.8 0.96 0.08
STG [38] 21.2 0.88 0.19 17.5 0.70 0.45 26.0 0.91 0.17
Method Single flag Tube Average
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
D-3DGS [36] 21.6 0.81 0.14 32.6 0.98 0.05 24.2 0.89 0.14
4DGS [37] 19.7 0.72 0.21 31.3 0.97 0.07 24.4 0.87 0.17
STG [38] 16.9 0.62 0.36 23.8 0.93 0.11 21.0 0.79 0.30
B.3 Additional Qualitative Results
Figure 10 shows additional qualitative results, where all methods fail to accurately capture multi-body interactions, producing blurred or under-reconstructed outputs. For the 4DGS [37], using the default grid learning rate frequently causes NaN values in the loss. Reducing the learning rate stabilizes training but results in poor dynamic capture, with outputs resembling static scenes.
(a) GT image
(b) D-3DGS [36]
(c) 4DGS [37]
(d) STG [38]
(e) SOM [15]
Figure 10: Qualitative results of recent DyNVS methods on the basin, bouncing balls, box-smoke, pancake, and ice scenes with monocular training setup. These results show that all methods frequently exhibit needle-like artifacts and fail to reconstruct dynamic elements accurately.
Appendix C License
Our dataset is released under the Creative Commons Attribution-NonCommercial (CC BY-NC) license. The mesh used for the lucy scene is sourced from the Stanford 3D Repository, which permits usage for research purposes. Other mesh objects are licensed under CC-BY 4.0, as detailed in Table 11. For the texture maps in the textile categories, we utilized images from Pixabay, which are freely available with contributor consent and comparable to a CC-BY 4.0 license, also summarized in Table 11. Therefore, all data included in our dataset comply with usage rights and do not pose intellectual property issues.
Table 11: License of the sources used for generating data.
        Scene         Resource name         License         Access
        Ice         Glass cup         CC BY 4.0         Sketchfab
        Hanok         Korean building         CC BY 4.0         Sketchfab
        Ship         Ship         CC BY 4.0         Sketchfab
        Pisa         Torre Pisa         CC BY 4.0         Sketchfab
        Bouncing balls         Fish bowl         CC BY 4.0         Sketchfab
        Cow         Cow         CC BY 4.0         Sketchfab
        Lucy         Lucy         research only         Stanford 3D Scan Repo
        Scene         Contributor         License         Access
        Flags         PaftDrunk         Content License         Pixabay
        Zoeysmom         Content License         Pixabay
        Single flag         –         Content License         Pixabay
        Lucy         WalterClark         Content License         Pixabay
        Tube         –         Content License         Pixabay
        Basin         Yourialka         Content License         Pixabay
Appendix D Limitations & Broad Impact
Limitations
Although our PhysGaia dataset holds a crucial position in the DyNVS field by enabling accurate evaluation with ground-truth physical properties, it is inherently limited to synthetic scenarios. To bridge this gap towards more realistic applications, we believe domain adaptation approaches, such as those from Syn2Real [79], could be adopted. However, such methods must be carefully designed to maintain multi-view consistency across frames. We leave this important domain transfer for future work, confident that our dataset’s precise physical information will be invaluable for developing these advanced techniques.
Broad impact
From a positive perspective, our work advances research on physically plausible 4D reconstruction. This, in turn, significantly improves monocular video reconstruction, an essential technology for future AR/VR applications. However, such progress in 4D reconstruction, especially given its potential for scene editing of existing videos, may raise intellectual property concerns regarding the original video content.

### v2

@@
-PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis
+PhysGaia: A Physics-Aware Benchmark with Multi-Body Interactions for Dynamic Novel View Synthesis
 Abstract
-We introduce PhysGaia, a novel physics-aware dataset specifically designed for Dynamic Novel View Synthesis (DyNVS), encompassing both structured objects and unstructured physical phenomena.
-Unlike existing datasets that primarily focus on photorealistic reconstruction, PhysGaia is created to actively support physics-aware dynamic scene modeling.
-Our dataset provides complex dynamic scenarios with rich interactions among multiple objects, where they realistically collide with each other and exchange forces.
-Furthermore, it contains a diverse range of physical materials, such as liquid, gas, viscoelastic substance, and textile, which moves beyond the rigid bodies prevalent in existing datasets.
-All scenes in PhysGaia are faithfully generated to strictly adhere to physical laws, leveraging carefully selected material-specific physics solvers.
-To enable quantitative evaluation of physical modeling, our dataset provides essential ground-truth information, including 3D particle trajectories and physics parameters, e.g., viscosity.
-To facilitate research adoption, we also provide essential integration pipelines for using state-of-the-art DyNVS models with our dataset and report their results.
-By addressing the critical lack of datasets for physics-aware modeling, PhysGaia will significantly advance research in dynamic view synthesis, physics-based scene understanding, and deep learning models integrated with physical simulation—ultimately enabling more faithful reconstruction and interpretation of complex dynamic scenes.
-Our datasets and codes are available in the project website [^1] http://cvlab.snu.ac.kr/research/PhysGaia.
+We introduce PhysGaia, a novel physics-aware benchmark for Dynamic Novel View Synthesis (DyNVS) that encompasses both structured objects and unstructured physical phenomena.
+While existing datasets primarily focus on photorealistic appearance, PhysGaia is specifically designed to support physics-consistent dynamic reconstruction.
+Our benchmark features complex scenarios with rich multi-body interactions, where objects realistically collide and exchange forces.
+Furthermore, it incorporates a diverse range of materials, including liquid, gas, textile, and rheological substance, moving beyond the rigid-body assumptions prevalent in prior work.
+To ensure physical fidelity, all scenes in PhysGaia are generated using material-specific physics solvers that strictly adhere to fundamental physical laws.
+We provide comprehensive ground-truth information, including 3D particle trajectories and physical parameters (e.g., viscosity), enabling the quantitative evaluation of physical modeling.
+To facilitate research adoption, we also provide integration pipelines for recent 4D Gaussian Splatting models along with our dataset and their results.
+By addressing the critical shortage of physics-aware benchmarks, PhysGaia can significantly advance research in dynamic view synthesis, physics-based scene understanding, and the integration of deep learning with physical simulation, ultimately enabling more faithful reconstruction and interpretation of complex dynamic scenes.
+Figure 1: Visualization of key properties of PhysGaia.
+Unlike existing benchmarks limited to single objects or materials, PhysGaia is a physics-aware benchmark featuring complex multi-body interactions across diverse substances (liquids, gases, rheological materials, and textiles).
+By providing ground-truth 3D trajectories and physical parameters, it uniquely enables the evaluation of physical realism alongside traditional photorealism.
+In addition to multi-body collisions, our dataset captures (c) splashing effects characterized by non-local rigid motion, as well as (d–e) complex optical phenomena such as specular reflection and refraction.
+footnotetext: indicates equal contribution.
 1 Introduction
-(a) Liquid with FLIP solver: hanok (left) and ship (right)
-(b) Gas with Pyro solver: box-smoke (left) and pisa (right)
-(c) Viscoelastic substances with MPM solver: pancake (left) and jelly party (right)
-(d) Textile with Vellum solver: lucy (left) and basin (right)
-Figure 1: Examples from the proposed physics-aware dataset, PhysGaia.
-They exhibit complex physical interactions between multiple objects composed of diverse materials such as liquid, gas, viscoelastic substance, and textile.
-This dataset will foster physics reasoning in dynamic scenes.
-Since the introduction of Neural Radiance Fields (NeRF) [1], the Novel View Synthesis (NVS) algorithms based on deep learning have progressed rapidly.
-While early research primarily targeted static scenes, recent efforts have extended the scope to dynamic scene reconstruction to better support diverse and interactive AR/VR applications.
-This emerging direction, known as Dynamic Novel View Synthesis (DyNVS), aims to reconstruct 4D scenes from input videos and synthesize photorealistic images at novel viewpoints and time steps that are not seen during training.
-The evolution of DyNVS has been closely tied to the development of suitable datasets.
-Early DyNVS datasets [2, 3, 4] primarily provided multiview training videos and typically involved only limited object motion.
-These datasets laid the foundation for early DyNVS research [3, 5, 6, 7, 8, 9, 10, 11], which focused on modeling scene dynamics by deforming canonical geometries over time.
-Recently, to better support AR/VR applications, datasets captured using handheld mobile devices have been introduced [7, 12, 13].
-These datasets consist of monocular videos, and recent research has primarily focused on mitigating overfitting under such limited supervision.
-Building on these recent datasets, subsequent works [14, 15, 16, 17, 18, 19, 20] also have commonly extended recent Gaussian Splatting [21] to dynamic scenarios, aiming to achieve photorealistic rendering.
-While these efforts have led to significant improvements in visual quality, photorealism alone is insufficient for presenting the underlying dynamics and physical properties of real-world scenes.
-As DyNVS continues to advance, a natural next step is to move beyond photorealism and incorporate physical realism—enabling models not only to render how scenes look, but also to reason about how they behave.
-Recent pioneering works [22, 23, 24, 25, 26] have begun to explore this direction by integrating differentiable physics simulation into Gaussian Splatting [21] frameworks.
-Despite this growing interest, current research largely remains limited to 4D generation [27, 28, 29, 30] or simplified DyNVS scenarios [26, 31], often restricted to single object or single material.
-Consequently, more complex scenarios—such as multi-object interactions or the modeling of diverse physical materials—are still significantly under-explored.
-In response, we introduce PhysGaia, a new benchmark designed to support and accelerate research in these emerging and challenging directions.
+Since the emergence of Neural Radiance Fields (NeRF), Novel View Synthesis (NVS) algorithms based on deep learning have witnessed rapid advancements.
+While initial research primarily focused on static environments, recent efforts have shifted toward dynamic scene understanding to meet the demands of immersive and interactive AR/VR applications.
+This emerging direction, termed Dynamic Novel View Synthesis (DyNVS), aims to reconstruct complex 4D spatiotemporal scenes from video inputs, enabling the synthesis of photorealistic images at novel viewpoints and time steps beyond the training distribution.
+The evolution of DyNVS has been closely related to the availability of diverse benchmarks.
+Early DyNVS datasets [40, 57, 80] primarily provided multi-view training videos, albeit with relatively constrained object motion.
+These benchmarks established the groundwork for initial research [57, 17, 12, 55, 8, 16, 61, 15], which focused on modeling temporal dynamics by deforming a canonical geometry over time.
+More recently, benchmarks featuring videos captured with handheld mobile devices have been introduced [55, 56, 18].
+These datasets consist of monocular sequences, prompting research to focus on mitigating overfitting under such sparse supervision.
+Building upon these foundations, subsequent works [33, 70, 84, 28, 36, 35, 7, 32] have extended Gaussian Splatting (GS) [31] to dynamic scenarios, aiming for high-fidelity, photorealistic rendering.
+As DyNVS continues to advance, a natural next step is to move beyond photorealism and incorporate physical realism, enabling models not only to render how scenes look, but also to reason about how they behave.
+Recent pioneering works [75, 83, 48, 4, 29] have begun exploring this direction by integrating differentiable physics simulation into the 4D Gaussian Splatting (4DGS) framework.
+Despite this burgeoning interest, existing research remains largely confined to 4D generation or simplified DyNVS scenarios [29, 85], often restricted to isolated objects with a single material type.
+Consequently, more complex environments—characterized by multi-body interactions and a diverse range of physical materials such as liquids and gases—remain significantly underexplored.
+Furthermore, while real-world videos of such complex scenarios can be readily captured, they inherently lack the ground-truth physical information (e.g., 3D trajectories and material parameters) necessary for a rigorous and quantitative evaluation of physical realism.
+In response, we introduce PhysGaia, a new benchmark with novel evaluation metrics designed to support and accelerate research in these challenging directions.
+Figure 1 illustrates examples with complex physical interactions in our benchmark.
 Our contributions are summarized as follows:
-  • We introduce PhysGaia, a physics-aware dataset featuring rich interactions among multiple objects and encompassing a wide range of physical materials, including liquids, gases, textiles, and viscoelastic substances, as illustrated in Figure 1.
+  • We introduce PhysGaia, a physics-aware dataset featuring rich interactions among multiple objects and encompassing a wide range of physical materials, including liquids, gases, textiles, and rheological substances.
   • PhysGaia provides essential ground-truth information, such as 3D particle trajectories and physical parameters, enabling quantitative evaluation of physical modeling.
-  • We evaluate existing state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and demonstrating the potential for improvement in this field.
-The rest of this paper is organized as follows.
-Section 2 reviews related work and highlights the uniqueness of our dataset.
-We discuss the main properties and potential research enabled by our dataset in Section 3.
-Section 4 details material-specific physics solvers used for dataset construction and methodologies for generating multi-body interaction scenarios.
-Section 5 shows our analysis, including evaluations of recent DyNVS algorithms on PhysGaia, We conclude the paper in Section 6.
+  • We test existing DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and demonstrating the potential for improvement in this field.
 2 Related Work
 2.1 Dynamic Novel View Synthesis
-In recent years, significant advances have been made in novel view synthesis [1, 32, 33, 34, 35, 21].
-Although this field initially have focused on static scene reconstruction but now it extended to handling dynamic scenes, which is known as Dynamic Novel View Synthesis (DyNVS).
-Early DyNVS methods were built on Neural Radiance Fields (NeRF) [5, 6, 7, 8, 9, 10, 11], which usually modeled scene dynamics either by implictly modeling with temporal inputs [5, 6] or directly estimating the time-wise deformation of canonical geometry through auxiliary neural networks [7, 8, 9, 10, 11].
-Following the emergence of 3D Gaussian Splatting (3DGS) [21], recent DyNVS research [17, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] has shifted toward Gaussian Splatting-based representations to achieve higher fidelity and real-time inference.
-Since Gaussian Splatting is an explicit representation, additional deformation networks can directly modify the scale, rotation, and position of each gaussian primitive over time.
-In a significant development, research is now moving towards constraining the deformation of gaussian primitives to adhere to physical realism.
-PhysGaussian [22] pioneered this by combining an MPM (Material Point Method) [48] simulator with Gaussian Splatting [21], treating each gaussian primitive as a particle within the MPM simulation.
-This work has inspired many subsequent studies [31, 26, 27, 28, 29, 30] integrating physics-aware priors into Gaussian Splatting; however, these efforts remain largely confined to generation tasks [27, 28, 29, 30], with only a few addressing Dynamic Novel View Synthesis (DyNVS) [31, 26].
-Even for DyNVS, methods are typically limited to single objects or single materials, with most focusing on viscoelastic substances.
-Consequently, exploring physics-aware DyNVS research involving rich object interactions and diverse physical materials remains significantly underexplored.
-We believe our dataset can serve as a foundational resource for this crucial direction.
+In recent years, significant advances have been made in novel view synthesis [52, 34, 9, 20, 69, 54, 31].
+Although this field initially focused on static scene reconstruction, it has since extended to dynamic scenes, known as Dynamic Novel View Synthesis (DyNVS).
+Early DyNVS methods were built on Neural Radiance Fields (NeRF) [17, 12, 55, 8, 16, 61, 15], which usually modeled scene dynamics either by implicitly modeling with temporal inputs [17, 12] or directly estimating the time-wise deformation of canonical geometry through auxiliary neural networks [55, 8, 16, 61, 15].
+Following the emergence of 3D Gaussian Splatting (3DGS)[31], recent DyNVS research has shifted toward Gaussian-based representations, leading to the development of 4D Gaussian Splatting (4DGS) [28, 79, 74, 42, 45, 49, 22, 43, 38, 13, 68, 47, 65, 32].
+In 4DGS, an additional deformation network is employed to animate canonical Gaussian primitives over time, enabling efficient and high-quality modeling of dynamic scenes.
+Modeling dynamic scenes with 4DGS is increasingly moving toward incorporating physical laws to govern motion.
+PhysGaussian [75] pioneers this direction by combining an MPM [66] simulator with Gaussian Splatting, treating each Gaussian primitive as a particle within the particle-grid simulation framework of MPM.
+This work has inspired many subsequent studies [85, 29, 30, 46, 27, 58] integrating physics-aware priors into 4DGS; however, these efforts remain largely confined to generation tasks [30, 46, 27, 58], with only a few addressing Dynamic Novel View Synthesis (DyNVS) [85, 29].
+Even among DyNVS methods, existing approaches are typically limited to single-object scenes with homogeneous materials, predominantly focusing on rheological substances.
+Consequently, physics-aware DyNVS involving rich object interactions and diverse physical materials remains largely underexplored, a gap our dataset is designed to address.
 2.2 4D Datasets for Dynamic Novel View Synthesis
-Table 1: Comparison with Dynamic Novel View Synthesis datasets.
-This table summarizes key attributes including the number of total scenes having test videos for DyNVS evaluation, the number of scenes with physical phenomena, presence of multi-body interaction, ground-truth physics parameters, e.g., viscosity, and ground-truth 3D particle trajectories.
-Our PhysGaia dataset uniquely offers a comprehensive set of diverse scenes with complex multi-body interactions and complete ground-truth physical information, a significant advancement over prior benchmarks.
-Datasets Scene Stats Physics Information Capture Type
-# total scene # physics Interaction Physics param. 3D Traj.
-Plenoptic [2]   6       1(gas) No No No Multiview
-D-NeRF [3]   8       1(viscoelastic) No No No Monocular
-NVIDIA Dynamic [4]   8       0 No No No Monocular
-Nerfies [7]   4       0 No No No Monocular
-HyperNeRF [12]   4       0 No No No Monocular
-DyCheck [13]   7       1(textile) No No No Monocular
-NeRF-DS [49]   8       0 No No No Monocular
-EvDNeRF [50]   6       0 No No No Multiview
-Synthetic Soccer [51]   3       0 No No No Multiview
-HDR-HexPlane [52]   8       0 No No No Multiview
-[…]
-Table 1 provides a comparative overview of our dataset alongside existing multiview and monocular DyNVS datasets.
-The initial DyNVS datasets [2, 3, 4] primarily employed multiview configurations or captured scenes with very limited motion, typically involving mostly rigid objects.
-For example, Plenoptic [2] was captured with sufficiently many multiview cameras at fixed positions.
-To advance DyNVS toward practical AR/VR applications, more user-friendly datasets, often captured using handheld mobile phones, were later introduced.
+Table 1: Comparison with Dynamic NVS datasets.
+Our PhysGaia benchmark offers diverse scenes with complex multi-body interactions and complete ground-truth physical information (parameters and trajectories).
+Datasets Scenes Physics Info. View
+Total Phys. Inter. Param. Traj.
+Plenoptic [40] 6 1(G) Multi
+NVIDIA Dynamic [80] 8 0 Mono
+FluidNexus [19] 240 240(G) Multi
+DyCheck [18] 7 1(T) Mono
+Nerfies [55] 4 0 Mono
+HyperNeRF [56] 4 0 Mono
+NeRF-DS [76] 8 0 Mono
+EvDNeRF [2] 6 0 Multi
+Synthetic Soccer [39] 3 0 Multi
+HDR-HexPlane [73] 8 0 Multi
+D-NeRF [57] 8 1(R) Mono
+Phystwin [29] 22 22(R) Multi
+PhysGaia (Ours) 17 17 Both
+G: Gas, T: Textile, R: Rheological
+Table 1 provides a comparative overview of our dataset alongside existing DyNVS datasets.
+The early DyNVS datasets [40, 57, 80] primarily employed multiview configurations or captured scenes with very limited motion, typically involving mostly rigid objects.
+These datasets paved the way for DyNVS research, with subsequent works [57, 17, 12, 55, 8, 16, 61, 15] exploring deforming canonical geometries over time to model scene dynamics.
+To advance DyNVS toward practical AR/VR applications, more user-friendly datasets captured using handheld mobile phones were later introduced.
 Nerfies [55] pioneered handheld iPhone captures, though its scenes remained largely static.
-HyperNeRF [12] then introduced more rapid and varied motions, and DyCheck [13] further resolved camera teleportation issues seen in HyperNeRF.
-These monocular datasets have motivated research mitigating overfitting to training videos in these settings.
-Some approaches [14, 15, 16] focus on leveraging additional priors such as diffusion models [53], depth estimation models [54], and point trackers [55], while others [15, 17, 18, 19, 20] emphasize constraining deformation by adopting techniques like motion factorization [56] or As-Rigid-As-Possible (ARAP) regularization [57].
-Nonetheless, the primary objective across these datasets remains photorealistic reconstruction, with limited emphasis on physics-aware dynamic modeling  [^2] Note that some recent datasets have also been introduced, but they are typically tailored to specific scenarios, such as event cameras [50], HDR rendering [52], specular lighting effects [49], or human motion capture [51]..
-Unlike existing DyNVS datasets, our PhysGaia dataset offers scenes with multi-body interactions involving diverse physical materials, as detailed in Table 1.
-While some datasets [2, 3, 13] include limited scenes with physical phenomena, these are typically few in number, not generated using accurate physics solvers, and lack complex multi-object interactions.
-In this context, PhysGaia occupies a unique position and holds strong potential to spur advancements in physics-aware dynamic scene modeling and reasoning.
+HyperNeRF [56] then introduced more rapid and varied motions, and DyCheck [18] further addressed camera teleportation artifacts observed in HyperNeRF.
+These monocular datasets have motivated research aimed at mitigating overfitting to training views, with some approaches [33, 84, 70] leveraging additional priors such as diffusion models [60], depth estimators [77], and point trackers [78], while others [28, 70, 35, 36, 7] constrain deformation via motion factorization [67] or As-Rigid-As-Possible (ARAP) regularization [64].
+Nonetheless, the primary objective across these datasets remains photorealistic reconstruction, with limited emphasis on physics-aware dynamic modeling.
+Unlike existing DyNVS datasets, PhysGaia provides scenes featuring multi-body interactions across diverse physical materials, as shown in Table 1.
+Although some datasets [40, 57, 18, 19, 29] include scenes with physical phenomena, they are limited in scale, lack accurate physics simulation, or do not capture complex multi-object interactions.
+In this regard, PhysGaia occupies a unique position with strong potential to advance physics-aware dynamic scene modeling.
 2.3 4D Datasets from Physics Simulator
-Table 2: Comparison with physics-simulated datasets.
-“Multi-body Inter.” denotes multi-body interaction.
-Unlike existing datasets, our PhysGaia offers rich interactions, a diverse range of physical materials, and access to accurate physics parameters and ground-truth 3D particle trajectories, making it a uniquely valuable resource for advancing physics-based understanding of dynamic scenes.
-Datasets Multi-body Inter. Materials Physics Information
-Liquids Gas Textile Viscoelastic Physics params. 3D traj.
+Table 2: Comparison with physics-simulated datasets for the DyNVS task.
+Our PhysGaia benchmark offers diverse materials, multi-body interactions, and complete physical ground truth (parameters and trajectories).
+Datasets Multi-body Materials Physics Info.
+Interactions L G T R Params. 3D Traj.
 DNG [81] Yes $\checkmark$ No Yes
@@
 Deng et al. [11] No $\checkmark$ No No
+ScalarFlow [14] No $\checkmark$ Yes Yes
 PAC-NeRF [41] No $\checkmark$ $\checkmark$ Yes Yes
 Spring-Gaus [85] No $\checkmark$ Yes Yes
-ScalarFlow [64] No $\checkmark$ Yes Yes
-[…]
-(a) PAC-NeRF [63]
-(b) Spring-Gaus [31]
-(c) ScalarFlow [64]
-Figure 2: Visualization of datasets most similar to our PhysGaia.
-While all of these datasets address physical phenomena, they are limited in several key aspects: limited coverage of physical materials, overly simplified dynamics, and an absence of rich multi-body interactions.
-To provide a comprehensive comparison, we evaluate our PhysGaia against other 4D datasets generated via physics simulation, as summarized in Table 2.
-Since DyNVS tasks require multiview RGB imagery, we focus on datasets that offer such data.
-While some existing datasets support multi-body interactions, they often lack rich interaction dynamics or are confined to a single material type.
-For instance, clothed human datasets [58, 59, 60] exhibit textile-human interactions but are limited to textiles and human-centric motions.
-In terms of ground-truth physical information, a few datasets [63, 31, 64] provide both physical parameters and 3D trajectories.
-However, these datasets are typically restricted to single materials and do not capture complex, multi-body interactions, as illustrated in Figure 2.
-Specifically, PAC-NeRF [63] (Figure 2(a)) focuses on liquids and viscoelastic materials, but its liquid scenarios are highly viscous, behaving similarly to viscoelastic materials.
-Spring-Gaus [31] (Figure 2(b)) is restricted to viscoelasticity, and ScalarFlow [64] (Figure 2(c)) is limited to non-interacting gases.
-In stark contrast, our PhysGaia directly addresses these limitations.
-It provides rich interactions, a diverse range of physical materials, and access to accurate simulation parameters and ground-truth trajectories.
-This makes PhysGaia a uniquely valuable dataset for advancing our understanding of physics in dynamic scenes.
-3 Dataset Properties and Research Impact
-We propose PhysGaia to advance physically realistic reconstruction in Dynamic Novel View Synthesis (DyNVS), extending the conventional reconstruction task beyond photorealism.
-As highlighted in Tables 1 and 2 of Section 2, our dataset uniquely features complex multi-body interactions involving a diverse range of physical materials, distinguishing it within the DyNVS landscape.
-We believe this comprehensive dataset holds significant potential to enhance the understanding of physics in dynamic scenes.
-This section outlines the key properties of our dataset and discusses potential future research directions it enables.
-3.1 Dataset Properties
-Complex physics-aware dynamics
-(a) Multi-body interaction: Jelly party
-(b) Multi-body interaction: Ship
-(c) Splashing (non-locally rigid)
-(d) Specular modeling
-(e) Refraction
-Figure 3: Visualization of physics properties in PhysGaia. Alongside multi-body interactions, PhysGaia also includes various physical phenomena like splashing, refraction, and specular effects.
-Our PhysGaia dataset comprises 17 scenes that include multi-body interactions.
-Figure 1 illustrates a subset of these scenes.
-We further visualize the detailed examples in Figure 3(a) and Figure 3(b), which exhibit collisions between viscoelastic objects causing rapid changes in motion direction, and rigid-liquid surface impacts resulting in splashing, respectively.
-To ensure that these scenes adhere to physical laws with accurately calculated force exchange among multi bodies, we carefully select material-specific solvers: FLIP for liquids, Pyro for gases, MPM for viscoelastic materials, and Vellum for textiles.
-Further details on simulation configurations for handling multi-body interactions are found in Section 4.
-Beyond multi-body interactions, our dataset also exhibits various physical phenomena such as non-locally rigid motion (common in liquid and gas scenes), specular reflection, and refraction, as shown in Figures 3(c)3(d), and 3(e), respectively.
-These properties enhance the realism of our dataset and enable a wide range of downstream tasks and research applications.
-Ground-truth physics information
-In contrast to real-world video datasets, which are easier to capture but lack access to underlying physical properties, our simulated dataset provides complete ground-truth physical information.
-This includes 3D particle trajectories and accurate physics parameters such as viscosity, Young’s modulus, Poisson’s ratio, and temperature for gas-based scenes [^3] Temperature is a crucial component for buoyancy calculation in smoke simulations..
-This comprehensive provision enables precise evaluation of physical reasoning in dynamic scenes and directly supports the future research directions discussed in Section 3.2.
+PhysGaia (Ours) Yes $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Yes Yes
+L: Liquid, G: Gas, T: Textile, R: Rheological
+To provide a comprehensive comparison, we also compare PhysGaia with existing 4D datasets generated via physics simulation, as summarized in Table 2.
+Since DyNVS tasks require multiview RGB imagery, we focus on datasets that provide such data.
+Some existing datasets support multi-object interactions, but either lack rich interaction dynamics or are limited to a single material type.
+For instance, clothed human datasets [81, 86, 71] naturally exhibit textile-body interactions, yet are restricted to textile materials and human-centric motions.
+In terms of ground-truth physical information, a few datasets [41, 85, 14] provide both physical parameters and 3D trajectories; however, they remain confined to single-material settings and do not capture rich multi-object interactions.
+PAC-NeRF [41] focuses on liquids and viscoelastic materials, but its liquid scenarios are constrained to highly viscous flows, making them behaviorally similar to viscoelastic materials.
+Spring-Gaus [85] is restricted to viscoelasticity, while ScalarFlow [14] concentrates on gas.
+In contrast, PhysGaia directly addresses these limitations by offering rich multi-object interactions, a diverse range of physical materials, and accurate simulation parameters and trajectories, making it a uniquely valuable resource for advancing physics-aware understanding of dynamic scenes.
+(a) Liquid with FLIP solver: hanok (left) and ship (right)
+(b) Gas with Pyro solver: box-smoke (left) and pisa (right)
+(c) Rheological substances with MPM solver: pancake (left) and jelly party (right)
+(d) Textile with Vellum solver: lucy (left) and basin (right)
+Figure 2: Examples from the proposed physics-aware benchmark, PhysGaia.
+They exhibit complex physical interactions between multiple objects composed of diverse materials such as liquid, gas, rheological substance, and textile.
+More importantly, our benchmark enables evaluation of physics realism to foster physics reasoning in dynamic scenes.
+3 PhysGaia
+We propose PhysGaia to advance physically realistic reconstruction in DyNVS, moving beyond mere photorealism.
+As shown in Tables 1 and 2, PhysGaia uniquely features complex multi-body interactions across a diverse range of physical materials, distinguishing it from existing DyNVS datasets.
+We believe PhysGaia holds significant potential to deepen the understanding of physics in dynamic scenes.
+3.1 Benchmark Properties
+Complex scenarios with physics-aware dynamics
+Our dataset consists of 17 scenes featuring multi-object interactions, as visualized in Figure 2.
+To ensure that each scene adheres to physical laws with accurately calculated force exchange among objects, we carefully select material-specific solvers: Fluid-Implicit Particle (FLIP) [5] for liquids, Pyro [62] for gases, Vellum [63] for textiles, and Material Point Method (MPM) [66] for rheological materials.
+Beyond multi-body interactions, our dataset also captures various physical phenomena such as non-locally rigid motion commonly observed in liquid and gas scenes, specular reflection, and refraction, as shown in Figure 1.
+These properties enhance the realism of our dataset and support a wide range of downstream tasks and research applications.
+Providing physics parameters
+In contrast to real-world video datasets, where the ease of capture is offset by inaccessible underlying physics, our simulated dataset offers complete access to all physical information.
+This includes 3D particle trajectories and physics parameters such as viscosity, Young’s modulus, Poisson’s ratio, and temperature for gas scenes [^1] Temperature is a crucial component when adopting smoke-related simulators for physical reasoning, particularly for buoyancy calculation..
+This comprehensive provision enables precise evaluation of physical reasoning in dynamic scenes, directly facilitating the future research directions outlined in Section 3.2.
 Supporting diverse DyNVS tasks
-Our dataset uniquely supports both multiview and monocular DyNVS task.
-Unlike most multiview datasets [2, 50, 51, 52] with videos captured from fixed camera positions like CCTV, PhysGaia provides several moving monocular video sequences with independent trajectories as training data.
+Our benchmark uniquely supports both multi-view and monocular dynamic novel view synthesis (DyNVS) tasks, as shown in Table 1.
+Unlike most existing multi-view datasets [40, 2, 39, 73] with videos captured from fixed camera positions, PhysGaia provides several moving monocular video sequences with independent trajectories.
 This allows for diverse training configurations; using full sequences corresponds to a multiview DyNVS task, while using a single sequence corresponds to a monocular DyNVS task.
-The monocular setup matches datasets like DyCheck [13], providing realistic handheld camera inputs without synthetic teleportation.
-For evaluation, we employ two static cameras with a large baseline.
+For the evaluation, we use two static cameras with a large baseline.
 Customizability
@@
 (d) Re-lighted
-Figure 4: Examples of diverse modalities that users can generate from the provided simulation node graphs.
-We provide the complete simulation node graphs and their exact parameter settings used to construct our dataset.
-These graphs encompass all relevant components, including physics solvers, source geometries, camera positions, lights, materials, and texture controls.
-By modifying these nodes, users can easily generate customized scenes or additional modalities such as depth maps, surface normals, and re-lighted images tailored to their specific downstream tasks, as shown in Figure 4.
+Figure 3: Examples of diverse modalities that users can generate from the provided simulation node graphs. This can facilitate adaptation to specific downstream tasks.
+We provide complete source files encompassing physics solvers, source geometries, materials, and texture controls.
+Such an open and comprehensive framework enables researchers to generate customized datasets with higher resolution imagery and diverse modalities including depth maps, surface normals, and relighted scenes, facilitating adaptation to specific downstream tasks, as shown in Figure 3.
 Accessibility
-Our PhysGaia benchmark is designed for research-friendly and easy access, providing integration pipelines that enable the use of state-of-the-art DyNVS models [15, 36, 37, 38] with our data, as detailed on our GitHub repository [^4] https://github.com/mjmjeong/PhysGaia.
+Our PhysGaia benchmark is designed for research-friendly and easy access, providing integration pipelines that enable the use of state-of-the-art 4D Gaussian Splatting models [70, 79, 74, 42, 38] with our data.
 Furthermore, we include COLMAP-reconstructed point clouds for each scene.
-These provisions aim to foster the adoption of our dataset by researchers working for state-of-the-art DyNVS models.
+These provisions aim to facilitate the adoption of our dataset by researchers working with state-of-the-art DyNVS models.
 3.2 Potential Research
-We highlight the potential impact of our PhysGaia dataset by outlining several promising future research directions it uniquely enables.
-Advancing physical reasoning in dynamic scenes
+This subsection highlights the potential impact of our PhysGaia dataset by outlining several promising future research directions it uniquely enables.
+Physical reasoning of dynamic scenes
 Since our PhysGaia dataset provides ground-truth physics information, it facilitates precise evaluation of physical reasoning in dynamic scenes.
 For instance, ground-truth physics parameters like viscosity can be used to evaluate inverse physics estimation methods, where differentiable simulators are employed to optimize these parameters.
-Furthermore, unlike existing DyNVS research that primarily focuses on photorealism and relies solely on ground-truth RGB images, our dataset offers ground-truth 3D trajectories.
-This enables evaluating the actual motion of individual Gaussian primitives in recent GS-based methods, moving beyond just visual fidelity.
-We believe this unique feature establishes our dataset as a valuable benchmark for developing and evaluating physics-aware DyNVS models./asset
-Exploring multi-body interactions
-While recent research integrates physics into DyNVS [26, 31], it largely remains limited to single materials and often single objects within scenes.
-As a result, the crucial aspect of physics reasoning for interactions between multiple objects in DyNVS—particularly for the estimation of force exchange and deformation during contact—remains largely unexplored.
-We believe our novel dataset—specifically designed to capture complex multi-body interactions—will serve as a valuable resource for advancing future research in multi-physics modeling and adaptive representations for hybrid scenes.
-Integrating material-specific physics solvers
-The dominant approach to incorporating physics into DyNVS algorithms currently relies on differentiable particle-based simulators [48, 65, 66, 67, 68, 69], which treat Gaussian primitives as particles.
-However, different physical phenomena are best captured by different physics solvers—a principle reflected in our dataset construction process described in Section 4.
-This insight can guide researchers seeking to integrate more appropriate solvers tailored to specific material behaviors, such as fluids, cloth, and smoke.
-For instance, the FLIP [70] solver excels in simulating incompressible fluids due to its hybrid particle-grid representation, offering improved stability and realism compared to purely particle-based SPH-based solvers.
-Similarly, thermodynamic phenomena, such as smoke dynamics, are conventionally handled by voxel-based simulators.
-Nevertheless, the seamless integration of these volumetric solvers into particle-based representations like Gaussian Splatting presents a significant and largely unexplored research challenge.
-4 Dataset Construction
-Our dataset consists of 17 scenes, categorized into four material types: liquids, gases, viscoelastic materials, and textiles.
-Each category contains 4 to 5 scenes.
+Furthermore, unlike existing 4DGS research that primarily focuses on photorealism and relies on ground-truth RGB images, our dataset offers ground-truth 3D trajectories, enabling evaluation of the actual motion of individual Gaussian primitives in 4DGS.
+We believe this unique feature establishes our dataset as a valuable benchmark for developing and evaluating physics-aware DyNVS models.
+Multi-body interaction
+While recent research integrates physics into DyNVS [29, 85, 6, 10], it largely remains limited to single materials and often single objects within scenes.
+As a result, the crucial aspect of physics reasoning for interactions between multiple objects in DyNVS – particularly the estimation of force exchange and deformation during contact – remains largely unexplored.
+We believe our novel dataset, specifically designed with complex multi-object interactions, will be instrumental in enabling significant future research in multi-physics modeling and adaptive representations for handling hybrid scenes.
+Integration of material-specific physics solver
+The dominant approach for integrating physics into DyNVS algorithms currently involves adopting differentiable simulators [66, 26, 51, 1, 24, 25], treating Gaussian primitives as particles in simulators.
+However, we emphasize that different physical phenomena are best captured by different physics solvers, a principle reflected in our dataset construction process detailed in our supplementary documents.
+This can guide researchers seeking to integrate more appropriate solvers tailored to specific material behaviors, such as fluids, cloth, or smoke.
+For example, the FLIP [5] solver excels at simulating incompressible fluids due to its hybrid particle-grid representation, offering greater stability and realism compared to purely particle-based SPH-based solver.
+Similarly, thermodynamic effects like temperature and buoyancy are crucial for smoke simulation, typically represented using voxel-based grids.
+Integrating such volumetric solvers into particle-based frameworks like Gaussian splatting, however, remains largely unexplored.
+4 Metrics for Physical Realism
+When reconstructing 4D scenes, faithfully capturing actual 3D particle flows can be viewed as a proxy for physically realistic reconstruction.
+To this end, we introduce metrics that quantify how well the motions of Gaussian primitives in 4DGS align with ground-truth physical trajectories.
+Trajectory Distance (TD)
+Trajectory distance measures the spatial deviation along entire trajectories.
+Let $M$ denote the number of reconstructed primitives, $T$ the total number of frames, $X_{i}^{t,\mathrm{recon}}$ the 3D position of reconstructed primitive $i$ at time $t$, and $X_{j}^{t,\mathrm{gt}}$ the 3D position of ground-truth trajectory $j$.
+TD is defined as the average Euclidean distance between matched trajectories as follows:
+$\mathrm{TD}=\frac{1}{MT}\sum_{i=0}^{M-1}\sum_{t=0}^{T-1}\|X_{i}^{t,\mathrm{recon}}-X_{j(i)}^{t,\mathrm{gt}}\|_{2},$ (1)
+where $j(i)$ denotes the closest ground-truth trajectory matched to primitive $i$ in the initial frame.
+Lower TD values indicate closer alignment with ground-truth motion, reflecting higher physical realism.
+Area Under the Outlier Percentages (AUOP)
+To complement TD, we introduce a percentage-based metric that captures outlier behavior holistically across the entire sequence.
+For each primitive $i$ at timestep $t$, an outlier indicator $O_{i}^{t}$ is defined as follows:
+$O_{i}^{t}=\begin{cases}1&\text{if }O_{i}^{t-1}=1\text{ or }\|X_{i}^{t,\mathrm{recon}}-X_{j(i)}^{t,\mathrm{gt}}\|_{2}>\delta,\\ 0&\text{otherwise},\end{cases}$ (2)
+where $1$ indicates the primitive is marked as an outlier, and $\delta$ is a deviation threshold.
+Once a trajectory exceeds this threshold, it remains marked as an outlier for all subsequent frames.
+The percentage of outlier trajectories is computed at each timestep, yielding a curve over time as shown in Figure 8.
+To summarize this temporal behavior, we propose AUOP, which quantifies the area under this curve and provides a single comprehensive value for the entire sequence.
+Lower AUOP values indicate better compliance with ground-truth physics, reflecting fewer and shorter deviations throughout the sequence.
+Discussion
+TD and AUOP provide complementary perspectives: TD measures absolute trajectory accuracy, while AUOP captures the persistence of large deviations, together enabling a more robust evaluation of physical realism in reconstructed flows.
+5 Empirical Analysis with PhysGaia
+In this section, we demonstrate the characteristics of our benchmark and highlight critical yet largely overlooked limitations of existing DyNVS approaches.
+5.1 Comparison with Physics-based Benchmarks
+Table 3: Comparison with existing physics-based benchmarks on motion complexity and photorealism.
+PhysGaia achieves the highest dynamic score [44] and lowest FID [23] and KID [3], demonstrating richer dynamics and greater visual realism than existing DyNVS datasets.
+Note that PAC-NeRF sequences are too short to compute a dynamic score.
+Benchmark Motion complexity Photorealism
+Multi-body Dynamic score $\uparrow$ FID $\downarrow$ KID $\downarrow$
+ScalarFlow [14] No 0.391 293.5 0.255
+PAC-NeRF [41] No N/A 242.6 0.164
+Spring-Gaus [85] No 0.372 261.8 0.171
+PhysGaia Yes 0.444 207.8 0.118
+(a) ScalarFlow [14]
+(b) PAC-NeRF [41]
+(c) Spring-Gaus [85]
+Figure 4: Limitations of existing datasets.
+While ScalarFlow, PAC-NeRF, and Spring-Gaus address physical phenomena, they are limited in narrow coverage of physical materials, overly simplified dynamics, and an absence of rich multi-object interactions.
+Existing physics-based benchmarks [14, 41, 85] [^2] These datasets provide reliable ground-truth physics, including trajectories and physical parameters, as detailed in Table 2.
+often exhibit limited motion diversity and simplistic dynamics, as visualized in Figure 4.
+Table 3 demonstrates that PhysGaia significantly bridges the gap in both motion complexity and realism.
+To evaluate motion complexity, we employ the dynamic score [44], which quantifies motion through feature-level distances and inter-frame variance.
+PhysGaia achieves the highest score, reflecting its capacity to capture intricate multi-body interactions.
+Regarding photorealism, we compute FID [23] and KID [3] by flattening video sequences into image sets to measure distributional alignment with TinyImageNet.
+Our dataset yields the lowest scores across both metrics, confirming superior visual fidelity.
+These results validate that PhysGaia effectively bridges the gap in existing benchmarks by integrating physical accuracy, high photorealism, and complex motion dynamics.
+5.2 Analysis on Dynamic Novel View Synthesis
+We analyze recent DyNVS approaches (D-3DGS [79], 4DGS [74], STG [42], MoSca [38], and SoM [70]) on PhysGaia in terms of both photorealism and physical realism.
+5.2.1 Evaluation of Photorealism
+Table 4: Average quantitative results for each material category across all algorithms.
+While performance is generally high for textile, it deteriorates significantly for rheological substances, which typically exhibit complex dynamics involving multiple interacting components.
+View Method Liquid Gas Rheological materials Textile
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+Monocular D-3DGS [79] 22.7 0.87 0.22 21.9 0.89 0.16 20.1 0.84 0.15 22.1 0.83 0.18
+4DGS [74] 24.2 0.87 0.23 21.7 0.88 0.17 19.5 0.82 0.18 24.9 0.84 0.18
+STG [42] 19.2 0.72 0.39 21.9 0.85 0.24 13.6 0.63 0.40 21.9 0.84 0.21
+MoSca [38] 20.5 0.78 0.39 21.2 0.79 0.35 17.8 0.74 0.23 18.6 0.74 0.36
+SoM [70] 19.6 0.80 0.32 20.0 0.84 0.27 16.7 0.75 0.23 20.7 0.79 0.22
+Multiview D-3DGS [79] 22.2 0.87 0.24 23.7 0.91 0.13 22.2 0.89 0.10 27.7 0.90 0.12
+4DGS [74] 25.1 0.88 0.22 24.2 0.89 0.17 21.0 0.85 0.15 26.6 0.87 0.15
+STG [42] 20.8 0.75 0.40 25.0 0.91 0.19 17.2 0.70 0.36 21.1 0.81 0.25
+(a) GT image
+(b) D-3DGS [79]
+(c) 4DGS [74]
+(d) STG [42]
+(e) MoSca [38]
+(f) SoM [70]
+Figure 5: Qualitative results of recent DyNVS methods on the jelly party scene with monocular setup.
+All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
+Please refer to our supplementary documents for more qualitative results.
+Figure 6: Comparison of motion diversity across DyNVS datasets.
+Dynamic Score (DS) [44] measures feature-level distances across frames, serving as a proxy for motion complexity.
+PhysGaia achieves the highest score, reflecting the rich multi-object interactions present in our benchmark.
+We adopt standard image quality metrics for photorealism evaluation: PSNR, SSIM [72], and LPIPS [82].
+Tables 4 and E show that existing methods struggle on PhysGaia, unlike their performance on existing DyNVS benchmarks [80, 18].
+We attribute this to the motion complexity from multi-object interactions, evidenced by the significantly higher motion entropy of PhysGaia (Figure 6).
+Each method faces distinct challenges: MLP- and grid-based representations (D-NeRF [57], 4DGS [74]) fail on fine-grained dynamics such as fluid splashing, polynomial motion models (STG [42]) are too restrictive for complex trajectories, and motion factorization or ARAP-based constraints (MoSca [38], SoM [70]) break down under rich multi-object interactions.
+Figure E illustrates these failures, where all methods produce needle-like artifacts in the jelly party scene.
+Additional qualitative results are provided in the supplementary material.
+Table 5: Average quantitative results for both monocular and multiview settings, averaged across all 17 scenes.
+While multiview setups generally offer better reconstruction performance than monocular ones, even multiview results achieve PSNR scores below 30.
+This highlights the substantial difficulty in reconstructing the complex multi-body interactions in our benchmark.
+Method Monocular Multiview
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+D-3DGS [79] 21.7 0.86 0.18 24.2 0.89 0.14
+4DGS [74] 22.7 0.85 0.19 24.4 0.87 0.17
+STG [42] 19.3 0.76 0.30 21.0 0.79 0.30
+MoSca  [38] 19.5 0.76 0.34 N/A N/A N/A
+SoM [70] 19.3 0.80 0.26 N/A N/A N/A
+5.2.2 Evaluation of Physical Realism
+(a) GT motion
+(b) D-3DGS [79]
+(c) 4DGS [74]
+(d) STG [42]
+(e) MoSca [38]
+(f) SoM [70]
+Figure 7: Comparison of reconstructed flows and ground truth on the Box-smoke and Pisa scenes.
+Reconstructed flows deviate significantly from ground truth, where photorealism is achieved through local surface fluctuations rather than following actual physical motion.
+Figure 8: Outlier percentage over time on the box-smoke scene with AUOP values indicated.
+The oracle (GT-GT) achieves near-zero outliers, while existing methods exhibit rapidly increasing outlier percentages, resulting in significantly higher AUOP values.
+This highlights their inability to track physically plausible flows.
+Table 6: Performance for physics realism. The box-smoke scene shows higher TD and AUOP; existing methods are hard to capture physics with simplified visuals.
+Method Cow Box-smoke
+TD $\downarrow$ AUOP $\downarrow$ TD $\downarrow$ AUOP $\downarrow$
+D-3DGS [79] 0.07   0.0 4.01 37.9
+4DGS [74] 0.90   0.0 6.00 68.8
+STG [42] 0.30   0.7 3.62 30.2
+MoSca [38] 0.33   0.7 3.67 28.3
+SoM [70] 0.61 17.7 4.45 25.4
+Oracle 0.03   0.0 0.84   0.3
+To evaluate whether Gaussian primitives follow ground-truth physical flows, we report TD and AUOP for two scenes in Table 6333Oracle denotes results from an additional set of ground-truth trajectories (GT-GT), serving as an upper bound on performance.: the less dynamic cow scene and the highly dynamic box-smoke scene.
+While all methods perform reasonably on the cow scene, they exhibit substantially higher TD and AUOP on the box-smoke scene—optimizing solely for RGB reconstruction achieves photorealism but fails to ensure physical plausibility.
+Figures 7 and 8 further confirm this: reconstructed trajectories deviate significantly from ground truth, with primitives merely fluctuating near the surface rather than following actual physical motion, and outlier percentages increasing rapidly over time.
+These results underscore the necessity of physics-aware evaluation and position PhysGaia as a critical benchmark for driving progress in this direction.
+5.3 Analysis on Inverse-physics Problems
+Table 7: Comparison of estimated physics parameters (Young’s modulus $E$ and Poisson’s ratio $\nu$) against ground truth.
+Both PAC-NeRF [41] and GIC [6] methods substantially underestimate material stiffness in multi-object scenarios.
+   Method    $E$    $\nu$
+   Ground Truth    $100.0\times 10^{3}$    0.50
+   PAC-NeRF [41]    $1.7\times 10^{3}$    0.42
+   GIC [6]    $23.4\times 10^{3}$    0.37
+Beyond DyNVS tasks, PhysGaia also supports physics parameter estimation [6, 85, 10] since it is faithfully generated using physics simulators with known ground-truth parameters.
+We evaluate both PAC-NeRF [41] and GIC [6] on the bouncing balls scene.
+As shown in Table 7, both methods exhibit similar trends and significantly underestimate Young’s modulus $E$.
+This behavior arises because both approaches were designed for single-object scenarios: when applied to multi-object scenes with complex motions and interactions, the algorithms favor lower $E$ values, allowing greater deformation to fit the observations.
+These results demonstrate that our benchmark exposes underexplored challenges in multi-object physics parameter estimation, providing a valuable testbed for advancing this research direction.
+6 Conclusion
+We propose a novel physics-aware dataset, specifically designed to understand physics in dynamic scenes, particularly for DyNVS.
+Our benchmark captures complex multi-body interactions with a wide variety of materials.
+Each scene is faithfully generated using material-specific physics solvers, ensuring adherence to physical laws and providing rich ground-truth physics data including 3D flows and configurations.
+This ground-truth data uniquely enables the evaluation of physical reasoning.
+We also test several tasks on PhysGaia, revealing their fundamental limitations in achieving physical realism and handling multi-objects, which highlights significant potential for improvement.
+Acknowledgements
+This work was partly supported by Samsung Research, Samsung Electronics Co., Ltd., the National Research Foundation of Korea (NRF) grant [RS-2022-NR070855, Trustworthy Artificial Intelligence], and the Institute of Information & communications Technology Planning & Evaluation (IITP) grants [RS-2025-25442338, AI star Fellowship Support Program (Seoul National University); RS-2022-II220959 (No.2022-0-00959), (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making; No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] funded by the Korea government (MSIT).
+A Preliminary: 4D Gaussian Splatting
+A.1 Gaussian primitive
+3D Gaussian Splatting has recently achieved real-time rendering with state-of-the-art quality on static scenes.
+It adopts an explicit representation of the 3D scene through a collection of Gaussian ellipsoids $\Gamma=\left\{\gamma_{1},...,\gamma_{K}\right\}$.
+Each primitive $\gamma_{k}$ is defined by an unnormalized 3D Gaussian kernel $\mathcal{G}_{k}(\bm{x})$, parameterized by its mean $\bm{\mu}_{k}$ and covariance $\bm{\Sigma}_{k}$:
+$\mathcal{G}_{k}(\bm{x};\bm{\mu}_{k},\bm{\Sigma}_{k})\coloneqq\exp\left(-\frac{1}{2}(\bm{x}-\bm{\mu}_{k})^{\top}\bm{\Sigma}_{k}^{-1}(\bm{x}-\bm{\mu}_{k})\right),$ (A)
+where $\bm{\mu}_{k}\in\mathbb{R}^{3}$ denotes the primitive’s center position, $\bm{\Sigma}_{k}\in\mathbb{R}^{3\times 3}$ is an anisotropic covariance matrix, and $\bm{x}\in\mathbb{R}^{3}$ represents an arbitrary spatial coordinate.
+Since $\bm{\Sigma}_{k}$ must remain positive semi-definite during optimization—a constraint that is difficult to enforce directly—it is instead factorized into a rotation matrix $\bm{R}_{k}$ and a scaling matrix $\bm{S}_{k}$:
+$\bm{\Sigma}_{k}\coloneqq\bm{R}_{k}\bm{S}_{k}\bm{S}_{k}^{\top}\bm{R}_{k}^{\top}.$ (B)
+Beyond these Gaussian parameters $(\bm{\mu}_{k},\bm{R}_{k},\bm{S}_{k})$, each primitive also requires an opacity value $\alpha_{k}\in[0,1]$ and a feature vector $\bm{f}_{k}\in\mathbb{R}^{d}$, which typically encodes RGB colors or spherical harmonic (SH) coefficients.
+Thus, a Gaussian primitive is fully specified as $\gamma_{k}\coloneqq(\bm{\mu}_{k},\bm{R}_{k},\bm{S}_{k},\alpha_{k},\bm{f}_{k})$.
+A.2 Differentiable rasterization
+Before rendering the Gaussian primitives $\Gamma$ into the image space, each 3D Gaussian kernel, $\mathcal{G}_{k}(\bm{x};\bm{\mu}_{k},\bm{\Sigma}_{k})$, is first projected onto the 2D image space, resulting in a 2D Gaussian kernel, $\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})$.
+Here, $\pi:\mathbb{R}^{3}\to\mathbb{R}^{2}$ denotes the projection from world coordinates to image space.
+In the projected Gaussian representation, $\bm{r}\in\mathbb{R}^{2}$ indicates a pixel location in an image, and the 2D mean $\bm{\mu}_{k}^{\pi}\in\mathbb{R}^{2}$ and covariance $\bm{\Sigma}_{k}^{\pi}\in\mathbb{R}^{2\times 2}$ are given by
+$\bm{\mu}_{k}^{\pi}\coloneqq\pi(\bm{\mu}_{k})\quad\quad\text{and}\quad\quad\bm{\Sigma}_{k}^{\pi}\coloneqq\bm{J}\bm{W}\bm{\Sigma}_{k}\bm{W}^{\top}\bm{J}^{\top},$ (C)
+where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation and $\bm{W}$ is the world-to-camera transformation matrix.
+When rendering the primitives in $\Gamma$ from a target camera view, they are sorted according to their depth relative to the camera center.
+The color of a pixel $\bm{r}$ is then determined by $\alpha$-blending:
+$\mathbf{\hat{I}}(\bm{r})\coloneqq\sum_{k=1}^{K}\omega^{\pi}_{k}(\bm{r})c(\bm{f}_{k},\bm{r}),$ (D)
+where $\omega^{\pi}_{k}(\bm{r})$ denotes the contribution of primitive $\gamma_{k}$ to pixel $\bm{r}$, and $c(\bm{f}_{k},\bm{r})$ is its corresponding color.
+If $\bm{f}_{k}$ encodes spherical harmonics coefficients, the color is evaluated using the view direction associated with $\bm{r}$; otherwise, $\bm{f}_{k}$ can directly represent the primitive’s RGB values.
+Following the $\alpha$-blending formulation in 3DGS [31], the visibility weight is defined as
+$\omega^{\pi}_{k}(\bm{r})\coloneqq\alpha_{k}\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})\prod_{j=1}^{k-1}\left(1-\alpha_{j}\mathcal{G}^{\pi}_{j}(\bm{r};\bm{\mu}_{j}^{\pi},\bm{\Sigma}_{j}^{\pi})\right),$ (E)
+where $\alpha_{k}\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})$ represents the opacity of the $k^{\text{th}}$ projected primitive at pixel $\bm{r}$, and the product term encodes the transmittance, i.e., how much light passes through the preceding primitives along the ray.
+Further details can be found in the original Gaussian Splatting paper [31].
+A.3 Deformation for Dynamic Modelings
+To represent 4D scenes with Gaussian Splatting, recent methods [79, 28, 74, 42, 70, 38] learn time-dependent deformations that map canonical 3D Gaussian primitives to their posed states over time.
+At time $t$, the position, rotation, and scale of the $k$-th primitive are updated as
+$(\bm{\mu}_{k}^{t},\bm{R}_{k}^{t},\bm{S}_{k}^{t})=\big(\bm{\mu}_{k}+\phi_{\mu}(\cdot,t),\;\phi_{r}(\cdot,t)\,\bm{R}_{k},\;\bm{S}_{k}+\phi_{s}(\cdot,t)\big),$ (F)
+where the deformation functions $\phi_{\mu}(\cdot,t)$, $\phi_{r}(\cdot,t)$, and $\phi_{s}(\cdot,t)$ denote learnable fields that may take as input various combinations of canonical primitive attributes, spatial coordinates, or time embeddings, depending on the parameterization.
+The rotation update $\phi_{r}(\cdot,t)$ is implemented as a residual rotation, typically parameterized via an exponential map to ensure $\phi_{r}(\cdot,t)\in SO(3)$.
+These functions are commonly instantiated with MLPs, learnable control points [70, 38, 28], HexPlane [74], or low-order polynomials [42].
+The resulting time-varying primitive is then given by $\gamma_{k}(t)\coloneqq(\bm{\mu}^{t}_{k},\bm{R}^{t}_{k},\bm{S}^{t}_{k},\alpha_{k},\bm{f}_{k})$.
+B Scene Composition
+Our benchmark consists of 17 scenes, categorized into four material types: liquids, gases, rheological materials, and textiles.
+Each category contains 4 to 5 scenes, as listed in Table A.
 This section details the solver properties for each material and provides key simulation details regarding multi-body interactions.
 Please refer to our supplemental document for other simulation configurations.
-4.1 Common Setup
+Table A: List of scenes included in our PhysGaia benchmark. Our benchmark consists of 17 scenes, categorized into four material types: liquids, gases, rheological materials, and textiles.
+Liquid Gas Rheological substance Textile
+Scene name Cereal Pisa Jelly party Lucy
+Ship Box-smoke Pancake Basin
+Hanok Single smoke Bouncing balls Flags
+Ice Falling Cow Single flag
+Tube
+Figure A and Figure B visualize more scenes in PhysGaia to supplement the Figure 3 in our main paper.
+As shown in Table 1 and Table 2 of the main paper, textile materials remain underexplored in physics-based datasets.
+To facilitate progressive research development in this domain, we include few simpler scenes such as single flag and tube that employ with only basic wind interactions.
+We believe this approach effectively supports advancements by providing more accessible starting points.
+Similarly, our benchmark also includes few simpler gas scenes, simple smoke and falling, for the same reason as with textile scenarios: to facilitate progressive development.
+For the cow scene, internal forces intrinsic to the cow itself introduce unique and more complex dynamics, rather than using only gravity.
+(a) Liquid: hanok (top-left), ship (top-right), cereal (bottom-left), ice (bottom-right)
+(b) Gas: box-smoke (top-left), pisa (top-right)), simple smoke (bottom-left), and falling (bottom-right)
+Figure A: Examples from the proposed physics-aware dataset, PhysGaia.
+They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, viscoelastic substance, and textile.
+This dataset will foster physics reasoning in dynamic scenes.
+(a) Rheological substances: pancake (top-left), jelly party (top-right), bouncing balls (bottom-left), and cow (bottom-right)
+(b) Textile: lucy (top-left), basin (top-right), flags (mid-left), tube (mid-right), and single flag (bottom)
+Figure B: Examples from the proposed physics-aware benchmark, PhysGaia.
+They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, rheological substance, and textile.
+This benchmark will foster physics reasoning in dynamic scenes.
+C Common Setup
 Simulator selection
@@
 All cameras follow spiral trajectories to capture diverse views over time.
-4.2 Liquid
-For liquid scenes, we adopt the Fluid-Implicit Particle (FLIP) solver [70], a hybrid particle-grid method.
+All simulations were conducted on 1 NVIDIA Titan Xp GPUs and Intel Xeon E5-2630 v4 CPUs with 40 cores.
+While most scenes rendered in 2–3 hours, the lucy and hanok scenes required 12–24 hours due to their complex geometry and increased simulation costs.
+Camera Trajectories
+(a) $180^{\circ}$ scene: 2 cams
+(b) $360^{\circ}$ scene: 4 cams
+Figure C: Visualization of training camera’s traj.
+All cameras follow spiral trajectories to capture diverse views over time.
+For the monocular setup, we select one of the training cameras to generate the training data.
+Figure C visualizes our training camera trajectories, where we adopt either 2 or 4 cameras depending on the properties of each scene.
+For the cow, single flag, and tube scenes–which resemble $180^{\circ}$ settings with background walls placed behind the objects–we use only 2 cameras, as shown in Figure 3(a).
+In contrast, for $360^{\circ}$ scenes, we place 4 cameras at evenly spaced viewpoints around the scene, enabling full coverage from four directions.
+All cameras follow spiral trajectories to capture diverse views over time.
+For the monocular setup, we select one of the training cameras to generate the training data.
+D Material-specific Physics Solver
+D.1 Liquid
+For liquid scenes, we adopt the Fluid-Implicit Particle (FLIP) solver [5] a hybrid particle-grid method.
 FLIP maintains particle velocities throughout the simulation, using the grid solely to compute and apply forces such as pressure and viscosity.
@@
 In contrast, for scenes like ship and cereal where objects fall into liquid, causing both fluid and objects to move and influence each other through force exchange, we use a dynamic operator to accurately model these more complex interactions.
-4.3 Gas
+Physics law
+We assume incompressible fluids in our simulations.
+This assumption simplifies the governing equations while maintaining high fidelity for typical liquid behaviors.
+Fluid motion is governed by the conservation of momentum and mass.
+We consider gravity as the sole external force.
+The momentum equation for our fluid, describing the conservation of momentum, is given by
+$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}\right)=-\nabla p+\mu\nabla^{2}\mathbf{u}+\rho\mathbf{g},$ (G)
+where $\rho$ is density, $\mathbf{u}$ is velocity, $t$ is time, $p$ is pressure, $\mu$ is dynamic viscosity, and $\mathbf{g}$ is gravitational acceleration.
+The incompressibility condition ensures the conservation of mass, stating that the fluid’s volume remains constant as follows:
+$\nabla\cdot\mathbf{u}=0,$ (H)
+where $\mathbf{u}$ is velocity.
+Physics parameters
+Table B summarizes the density ($\rho$) and viscosity ($\mu$) values, which are the most important factors governing fluid motion.
+For the ship scene, we reduced the particle separation from 0.1 to 0.05 to decrease initial particle displacement and improve simulation accuracy.
+Also, the particle separation during dynamic simulation is set to 0.1 for the cereal scene and 0.05 for the ship scene.
+Table B: Physics parameters used for liquid materials.
+Scene name Object Density ($kg/m^{3}$) Viscosity
+Ship Water 800 0
+Ship 300 -
+Cereal Water 1000 0
+Cereal 1000 -
+Hanok Snow 1000 10
+Ice Water 1000 0
+D.2 Gas
 For gas (smoke) simulation, we utilize the Pyro solver [62].
@@
 Other scenes involving gaseous materials utilized the default simulation settings.
-4.4 Viscoelastic Substances
-We use MPM [48] for viscoelastic substances like snow and jelly.
+Physics law
+Our gas simulations, similar to liquids, are governed by the Navier-Stokes equations.
+However, we specifically account for the effects of buoyancy as an additional external force due to temperature differences.
+$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}\right)=-\nabla p+\mu\nabla^{2}\mathbf{u}+\rho(\mathbf{g}+\mathbf{f}_{b}),$ (I)
+where $\mathbf{f}_{b}$ is the buoyancy force. The buoyancy force is derived from temperature differences relative to the ambient environment as follows:
+$\mathbf{f}_{b}=k_{B}\cdot(T-T_{\text{ambient}})\cdot\mathbf{d}_{B},$ (J)
+where $k_{B}$ is the buoyancy constant, $T$ is the local gas temperature, $T_{\text{ambient}}$ is the ambient temperature, and $\mathbf{d}_{B}$ is the buoyancy direction.
+Physics parameters
+In gas simulations, temperature is the most critical physical parameter, as it directly governs buoyancy, expansion, and overall flow dynamics.
+We initialized the temperature of gas materials to 3000K.
+As temperature evolves locally due to gas movement and interactions, its spatiotemporal variation is represented as a temperature field defined on a voxel grid.
+These temperature fields are accessible via the provided simulation source file with a Python API, similar to how flow fields are accessed for generating ground-truth trajectory data.
+D.3 Rheological Substances
+We use MPM [66] for rheological substances like snow and jelly.
 As an extension of FLIP [5], MPM is ideal for simulating chunk-based materials.
@@
 Additionally, for the bouncing balls scene, we added a static bowl-shaped collider so that the falling balls rebound off both one another and the bowl’s surface.
-4.5 Textile
+Physics law
+Simulating viscoelastic materials requires modeling both their elastic (solid-like) and viscous (fluid-like) responses.
+Our approach follows fundamental conservation laws while adopting simplified constitutive models commonly used in MPM-based visual simulations, namely linear elasticity and Kelvin–Voigt viscosity.
+The conservation of mass is described by the continuity equation as follows:
+$\frac{D\rho}{Dt}+\rho(\nabla\cdot\mathbf{v})=0,$ (K)
+where $\rho$ is the material density and $\mathbf{v}$ is the velocity field. Conservation of linear momentum is given by
+$\rho\frac{D\mathbf{v}}{Dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{g},$ (L)
+where $\bm{\sigma}$ is the Cauchy stress tensor and $\mathbf{g}$ is gravitational acceleration.
+We adopt a Kelvin–Voigt-type viscoelastic model, where the total stress is decomposed into elastic and viscous components as follows:
+$\bm{\sigma}=\bm{\sigma}_{\text{elastic}}+\bm{\sigma}_{\text{viscous}}.$ (M)
+The elastic stress is modeled using linear isotropic elasticity as follows:
+$\bm{\sigma}_{\text{elastic}}=\frac{E\nu}{(1+\nu)(1-2\nu)}(\text{tr}(\bm{\epsilon}))\mathbf{I}+\frac{E}{2(1+\nu)}\bm{\epsilon},$ (N)
+where $E$ is Young’s modulus, $\nu$ is Poisson’s ratio, and $\bm{\epsilon}$ is the strain tensor defined as $\bm{\epsilon}=\frac{1}{2}(\nabla\mathbf{u}+(\nabla\mathbf{u})^{T})$, where $\mathbf{u}$ is the displacement field.
+The viscous stress accounts for rate-dependent deformation as follows:
+$\bm{\sigma}_{\text{viscous}}=2\mu\left(\mathbf{E}-\frac{1}{3}(\text{tr}(\mathbf{E}))\mathbf{I}\right)+\zeta(\text{tr}(\mathbf{E}))\mathbf{I},$ (O)
+where $\mu$ is the dynamic viscosity and $\zeta$ is the bulk viscosity. The strain-rate tensor $\mathbf{E}$ is defined as $\mathbf{E}=\frac{1}{2}(\nabla\mathbf{v}+(\nabla\mathbf{v})^{T})$.
+Physics parameters
+Gravity is applied to all scenes; however, for the cow scene, we introduce additional internal forces to induce cow motion and increase the complexity of the physical reasoning scenarios.
+Boundaries are set to be open in all directions, allowing objects to move freely without collisions against invisible walls.
+Table C lists detailed physics parameters, including Young’s modulus ($E$) and Poisson’s ratio ($\nu$), used in the simulations.
+Table C: Physics parameters used for rheological materials.
+Scene name Object Type $E$ $\nu$ Viscosity
+Pancake Honey Viscous $2.5\times 10^{5}$ 0.10 0.125
+Pancakes Chunky $8.0\times 10^{4}$ 0.23 -
+Bouncing balls Balls Elastic $1.0\times 10^{5}$ 0.50 -
+Fishbowl Static - - -
+Jelly party Jelly Elastic $8.0\times 10^{4}$ 0.45 -
+Cow Cow Elastic $5.0\times 10^{4}$ 0.45 -
+D.4 Textile
 For textile materials, we adopt the Vellum solver [63], which is based on the Extended Position Based Dynamics (XPBD) framework [50].
@@
 Dynamic interactions
-To simulate interactions between objects and textiles–where both move and exchange forces, as seen particularly in the basin scene–we employ the shape match constraint.
-This constraint helps maintain the overall shape of objects by driving points toward their rest configuration, allowing the material to preserve its structural integrity while still interacting dynamically with textiles and other objects.
+To simulate interactions between objects and textiles–where both move and exchange forces, as seen particularly in the basin scene–we employ the shape match constraint on objects.
+This constraint helps maintain the overall shape of rigid objects by driving points toward their rest configuration, allowing the material to preserve its structural integrity while still interacting dynamically with textiles and other objects.
 For the lucy scene, we increased the simulation sub-step count fivefold over the default to robustly handle collisions with the statue’s complex geometry.
 To simulate wind-induced fluttering in the flags, single-flag, and tube scenes, we applied external forces using a POP Wind node that blows parallel to the ground plane.
-5 Analysis
-5.1 Evaluation of Existing Algorithms on PhysGaia
-Experimental setup
-We evaluate recent DyNVS methods on our PhysGaia dataset: D-3DGS [36], 4DGS [37], STG [38], and SOM [15].
-Except for SOM [15], which is specialized for monocular setups, all models were tested under both monocular and multiview setups.
-For point cloud initialization, we first run COLMAP for dense matching and fusion, then uniformly downsample its output to approximately 40,000 points.
-All experiments are conducted on a NVIDIA A6000 GPU.
-For evaluation, we adopt standard image quality metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [77], and Learned Perceptual Image Patch Similarity (LPIPS) [78].
-Quantitative results
-Table 3: Average quantitative results for each material category across all baseline algorithms.
-Performance tends to be higher for textile materials, but it notably worsens for viscoelastic materials, which typically exhibit highly complex motions involving numerous independent objects.
-Capture Type Method Liquid Gas
-PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Monocular D-3DGS [36] 22.7 0.87 0.22 21.9 0.89 0.16
-4DGS [37] 24.2 0.87 0.23 21.7 0.88 0.17
-STG [38] 19.2 0.72 0.39 21.9 0.85 0.24
-SOM [15] 19.6 0.80 0.32 20.0 0.84 0.27
-Multiview D-3DGS [36] 22.2 0.87 0.24 23.7 0.91 0.13
-4DGS [37] 25.1 0.88 0.22 24.2 0.89 0.17
-STG [38] 20.8 0.75 0.40 25.0 0.91 0.19
-Capture Type Method Viscoelastic materials Textile
-PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Monocular D-3DGS [36] 20.1 0.84 0.15 22.1 0.83 0.18
-4DGS [37] 19.5 0.82 0.18 24.9 0.84 0.18
-STG [38] 13.6 0.63 0.40 21.9 0.84 0.21
-SOM [15] 16.7 0.75 0.23 20.7 0.79 0.22
-Multiview D-3DGS [36] 22.2 0.89 0.10 27.7 0.90 0.12
-4DGS [37] 21.0 0.85 0.15 26.6 0.87 0.15
-STG [38] 17.2 0.70 0.36 21.1 0.81 0.25
-Tables 3 and 4 report average performance by material type and across all 17 scenes.
-As expected, multiview setups outperform monocular ones, yet even they often yield PSNR below 30.
-This underscores the difficulty of capturing complex multi-body dynamics in our dataset, suggesting physics-aware priors are essential beyond RGB-based fitting alone.
-Among materials, textiles are generally reconstructed more accurately, while viscoelastic materials pose greater difficulty due to their complex, object-rich motion.
-Interestingly, SOM [15], despite strong results on other datasets, performs inconsistently on ours.
-We presume its motion factorization struggles with highly dynamic phenomena such as splashing and multiple interacting objects.
-Additionally, its heavy reliance on external modules like point trackers leads to performance drops when these modules fail in challenging scenes, particularly those lacking distinct visual features like smoke or gas.
-Qualitative results
-(a) GT image
-(b) D-3DGS [36]
-(c) 4DGS [37]
-(d) STG [38]
-(e) SOM [15]
-Figure 5: Qualitative results of recent DyNVS methods on the jelly party, lucy, and hanok scenes with monocular setup.
-All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
-Figure 5 presents qualitative results from monocular settings, revealing that all methods struggle to accurately capture multi-body interactions.
-This leads to common issues like needle-like artifacts and under-reconstruction of dynamic regions.
-For 4DGS [37], its default grid learning rate often causes NaN loss values; while reducing it stabilizes training, the output resembles static scenes due to poor dynamic capture.
-More results are available in the supplementary material.
-5.2 Limitation in Physical Realism
-We analyze a key limitation of recent DyNVS methods regarding physical realism.
-Since Gaussian Splatting-based representation, used in recent algorithms, is an explicit representation, the motion of gaussian primitives is expected to closely follow the actual physical trajectories of dynamic scenes.
-However, all tested DyNVS algorithms consistently fail to accurately reconstruct these physically reliable motions, as shown in Figure 6.
-D-3DGS [36], 4DGS [37], and STG [38] produce only localized fluctuations, while SOM [15] exhibits erratic motion primarily due to inaccuracies from external point trackers.
-This discrepancy is acutely apparent in fluid or gas scenes.
-Here, photorealism is easily achieved via color similarity alone, allowing primitives to simply fluctuate locally near the surface rather than genuinely modeling true dynamics.
-This analysis highlights a critical limitation—relying solely on RGB reconstruction loss ensures photorealism but does not guarantee physical realism.
-We believe that our PhysGaia dataset, which provides ground-truth trajectories, will be instrumental in advancing future research toward physics-aware dynamic scene reconstruction.
-Table 4: Average quantitative results across all 17 scenes for monocular and multiview settings.
-Multiview setups generally offer better reconstruction performance than monocular ones, yet PSNR scores remain below 30 even with multiple views.
-This highlights the inherent difficulty of reconstructing complex multi-body interactions within our dataset.
+Physics law
+We simulate textile materials using the Vellum solver [63], which is based on Extended Position-Based Dynamics (XPBD) [50].
+XPBD extends classical Position-Based Dynamics (PBD) by introducing a compliance parameter, allowing constraint stiffness to be decoupled from the time step and iteration count.
+Unlike force-based methods, XPBD does not explicitly solve equations of motion.
+Instead, particle positions are updated by iteratively projecting them to satisfy geometric constraints.
+The core algorithm for each simulation substep is shown in Algorithm A.
+Algorithm A Extended Position-Based Dynamics (XPBD) Algorithm (per substep)
+1:Input:
+[…]
+[…]
+[…]
+[…]
+6: Iterations: Number of constraint projection iterations
+7:// 1. Predict positions
+8:for each particle $i$ do
+9:  $\mathbf{p}^{*}_{i}\leftarrow\mathbf{p}_{i}+\Delta t\mathbf{v}_{i}+\Delta t^{2}\mathbf{a}_{\text{ext},i}$
+10:end for
+11:// 2. Constraint projection
+12:for $k=1$ to Iterations do
+13:  for each constraint $C_{j}$ do
+14:   compute $\Delta\mathbf{p}_{i}$ using Eq. (P)
+15:   $\mathbf{p}^{*}_{i}\leftarrow\mathbf{p}^{*}_{i}+\Delta\mathbf{p}_{i}$
+16:  end for
+17:end for
+18:// 3. Velocity update
+19:for each particle $i$ do
+20:  $\mathbf{v}_{i}\leftarrow(\mathbf{p}^{*}_{i}-\mathbf{p}_{i})/\Delta t$
+21:  $\mathbf{p}_{i}\leftarrow\mathbf{p}^{*}_{i}$
+22:end for
+We primarily employ stretch and bend constraints to model textile behavior.
+For a constraint $C(\mathbf{p})$, where $C(\mathbf{p})$ is a scalar constraint function that measures the violation of a geometric constraint, such as maintaining distances (stretch) or bending angles between particles, with $C(\mathbf{p})=0$ indicating that the constraint is satisfied, the position correction is as follows:
+$\Delta\mathbf{p}_{i}=-\Delta\lambda\,\mathbf{w}_{i}\nabla_{\mathbf{p}_{i}}C,$ (P)
+where
+$\Delta\lambda=\frac{C(\mathbf{p})}{\sum_{j}\mathbf{w}_{j}|\nabla_{\mathbf{p}_{j}}C|^{2}+\tilde{\alpha}},\quad\tilde{\alpha}=\frac{\alpha}{\Delta t^{2}}.$ (Q)
+Here, $\alpha$ denotes the compliance parameter (inverse stiffness).
+Physics parameters
+For textiles, we primarily adopt the default configuration from the Vellum solver, with detailed modifications summarized in Table D.
+In addition to stiffness, we specify a damping ratio for each constraint to reduce oscillations and stabilize the simulated cloth.
+Table D: Physics parameters used for textile materials.
+Scene name Object Stretch Bend
+Stiffness Damping Ratio Stiffness Damping Ratio
+Flags Flag1 (small) $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+Flag2 (big) $\times 10^{11} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+Basin Boxes $\times 10^{10} 1.0×10$$\times 10^{-3}$ N/A N/A
+Cloth $\times 10^{13} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+Single flag Flag $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+Lucy Cloth $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+Tube Tube flag $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
+E Details of Experimental Results
+E.1 Implementation details
+We implement recent DyNVS methods, including D-3DGS [79], 4DGS [74], STG [42], MoSca [38], and SOM [70].
+We follow the default training settings provided for each method.
+For 4DGS, we adopt the training configuration used for the HyperNeRF [56] dataset and reduce the grid learning rate to improve training stability.
+For SOM, we apply the recommended hyperparameters for the DyCheck [18]dataset.
+For point cloud initialization, we run COLMAP with dense matching and fusion, followed by uniform downsampling to approximately 40,000 points.
+All experiments are conducted on NVIDIA RTX A5000 and A6000 GPUs, with training times ranging from 30 minutes to 2 hours depending on the method.
+Note that all experiments are implemented based on their public codes  [^4] D-3DGS: https://github.com/ingra14m/Deformable-3D-Gaussians [^5] STG: https://github.com/oppo-us-research/SpacetimeGaussians6664DGS: https://github.com/hustvl/4DGaussians [^7] SOM: https://github.com/vye16/shape-of-motion [^8] MoSca: https://github.com/JiahuiLei/MoSca.
+E.2 Average performance: all scenes
+Table E presents the average performance of all methods under both monocular and multiview settings, aggregated across all 17 scenes.
+For SOM [70], we observed better performance when scaling the estimated depth map using the COLMAP point cloud obtained from dense matching, compared to sparse matching.
+Table E: Average quantitative results for both monocular and multiview settings, averaged across all 17 scenes.
+While multiview setups generally offer better reconstruction performance than monocular ones, even multiview results achieve PSNR scores below 30.
+This highlights the substantial difficulty in reconstructing the complex multi-body interactions in our benchmark.
 Method Monocular Multiview
@@
 STG [42] 19.3 0.76 0.30 21.0 0.79 0.30
+MoSca [38] 19.5 0.76 0.34 N/A N/A N/A
 SoM [70] 19.3 0.80 0.26 N/A N/A N/A
-(a) GT image
-(b) D-3DGS [36]
-(c) 4DGS [37]
-(d) STG [38]
-(e) SOM [15]
-Figure 6: Comparison of reconstructed trajectories and its ground truth on the Box-smoke and Pisa scenes. The reconstructed trajectories do not accurately follow the true physical trajectories.
-6 Conclusion
-We propose a novel physics-aware dataset, PhysGaia, specifically designed to understand physics in dynamic scenes, particularly for Dynamic Novel View Synthesis (DyNVS).
-Comprising 17 diverse scenes, our dataset captures complex multi-body interactions with a wide variety of materials.
-Each scene is faithfully generated using material-specific physics solvers, ensuring adherence to physical laws and providing rich ground-truth physics data including 3D particle trajectories and physics parameters.
-This ground truth data uniquely enables the evaluation of physics reasoning.
-We also test state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and highlighting significant potential for improvement.
-We believe PhysGaia will be a critical resource that accelerates progress in physics-aware dynamic scene understanding.
-Appendix A Details of PhysGaia
-A.1 Scene Lists
-Our dataset consists of 17 scenes divided into four categories: liquid, gas, viscoelastic substances, and textile.
-Each category contains 4 to 5 scenes, and the detailed scene names are listed in Table 5.
-Table 5: List of scenes included in our PhysGaia dataset
-Liquid Gas Viscoelastic substance Textile
-Scene name Cereal Pisa Jelly party Lucy
-Ship Box-smoke Pancake Basin
-Hanok Single smoke Bouncing balls Flags
-Ice Falling Cow Single flag
-Tube
-A.2 Camera Trajectories
-(a) $180^{\circ}$ scene: 2 cams
-(b) $360^{\circ}$ scene: 4 cams
-Figure 7: Visualization of training camera’s traj.
-Figure 7 visualizes our training camera trajectories, where we adopt either 2 or 4 cameras depending on the properties of each scene.
-For the cow, single flag, and tube scenes–which resemble $180^{\circ}$ settings with background walls placed behind the objects–we use only 2 cameras, as shown in Figure 7(a).
-In contrast, for $360^{\circ}$ scenes, we place 4 cameras at evenly spaced viewpoints around the scene, enabling full coverage from four directions.
-All cameras follow spiral trajectories to capture diverse views over time.
-For the monocular setup, we select one of the training cameras to generate the training data.
-A.3 Implementation Details
-All simulations were conducted on 4 NVIDIA Titan Xp GPUs and Intel Xeon E5-2630 v4 CPUs with 40 cores.
-While most scenes rendered in 2–3 hours, the lucy and hanok scenes required 12–24 hours due to their complex geometry and increased simulation costs.
-To supplement Section 4, we provide additional detailed simulation implementation along with the value of physics parameters.
-These parameters can be served as ground-truth labels for downstream physical reasoning tasks.
-Liquid
-We assume incompressible, non-Newtonian fluids in our simulations.
-Table 6 summarizes the density and viscosity values, which are the most important factors governing fluid motion.
-External forces are limited to gravity.
-For the ship scene, we reduced the particle separation from 0.1 to 0.05 to decrease initial particle displacement and improve simulation accuracy.
-Also, the particle separation during dynamic simulation is set to 0.1 for the cereal scene and 0.05 for the ship scene.
-Table 6: Physics parameters used for liquid materials.
-      Scene name       Object       Density ($kg/m^{3}$)       Viscosity
-      Ship       Water       800       0
-      Ship       300       -
-[…]
-      Cereal       1000       -
-[…]
-      Ice       Water       1000       0
-Gas
-In gas simulations, temperature is the most critical physical parameter, governing buoyancy, expansion, and flow dynamics.
-We initialized the temperature to 3000K for gas materials.
-As temperature evolves locally due to gas movement and interactions, its spatiotemporal variation is represented as a temperature field defined on a voxel grid.
-These temperature fields are accessible via the provided simulation source file with Python API, similar to how flow fields are accessed for generating ground-truth trajectory data.
-Viscoelastic substances
-Gravity is applied to all scenes; however, for the cow scene, we introduce additional internal forces to induce cow motion and increase the complexity of the physical reasoning scenarios.
-Boundaries are set to be open in all directions, allowing objects to move freely without collisions against invisible walls.
-Table 7 lists detailed physics parameters, including Young’s modulus ($E$) and Poisson’s ratio ($P$), used in the simulations.
-Table 7: Physics parameters used for viscoelastic materials.
-Scene name Object Type $E$ $P$ Viscosity
-Pancake Honey Viscous $2.5\times 10^{5}$ 0.10 0.125
-Pancakes Chunky $8.0\times 10^{4}$ 0.23 -
-[…]
-Fishbowl Static - - -
-[…]
-Cow Cow Elastic $5.0\times 10^{4}$ 0.45 -
-Textile
-For textiles, we primarily adopt the default configuration from the Vellum solver, with detailed modifications summarized in Table 8.
-Table 8: Physics parameters used for textile materials.
-Scene name Object Stretch Bend
-Stiffness Damping Ratio Stiffness Damping Ratio
-Flags Flag1 (small) $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
-Flag2 (big) $1.0\times 10^{11}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
-[…]
-Cloth $1.0\times 10^{13}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
-[…]
-Lucy Cloth $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
-Tube Tube flag $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
-A.4 Scene Visualization
-Figure 8 and Figure 9 visualize more scenes in PhysGaia to supplement the Figure 1 in our main paper.
-As shown in Table 2 of the main paper, textile materials remain underexplored in physics-based datasets.
-To facilitate progressive research development in this domain, we include few simpler scenes such as single flag and tube that employ with only basic wind interactions.
-We believe this approach effectively supports advancements by providing more accessible starting points.
-Similarly, our dataset also includes few simpler gas scenes, simple smoke and falling, for the same reason as with textile scenarios: to facilitate progressive development.
-For the cow scene, internal forces intrinsic to the cow itself introduce unique and more complex dynamics, rather than using only gravity.
-(a) Liquid: hanok (top-left), ship (top-right), cereal (bottom-left), ice (bottom-right)
-Figure 8: Examples from the proposed physics-aware dataset, PhysGaia.
-(a) Gas: box-smoke (top-left), pisa (top-right)), simple smoke (bottom-left), and falling (bottom-right)
-(b) Viscoelastic substances: pancake (top-left), jelly party (top-right), bouncing balls (bottom-left), and cow (bottom-right)
-(c) Textile: lucy (top-left), basin (top-right), flags (mid-left), tube (mid-right), and single flag (bottom)
-Figure 9: Examples from the proposed physics-aware dataset, PhysGaia.
-They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, viscoelastic substance, and textile.
-This dataset will foster physics reasoning in dynamic scenes.
-Appendix B Details of Experimental Results
-To supplement Table 3, Table 4, and Figure 5 of the main paper, we report more detailed quantitative results along with additional qualitative results.
-B.1 Performance breakdown: monocular setting
-We provide per-scene breakdown performance of monocular video reconstruction setting in Table 9.
-We report the performance of D-3DGS [36], 4DGS [37], STG [38], and SOM [15], which serve as the most common and recent baselines for the DyNVS task.
-Table 9: Per-scene breakdown results for all 17 scenes under the monocular setting.
+E.3 Performance breakdown: monocular setting
+We provide per-scene breakdown performance of monocular video reconstruction setting in Table F.
+We report the performance of D-3DGS [79], 4DGS [74], STG [42], MoSca [38] and SOM [70], which serve as the most common and recent baselines for the DyNVS task.
+Table F: Per-scene breakdown results for all 17 scenes under the monocular setting.
 Method Cereal Ship Hanok
@@
 STG [42] 15.8 0.54 0.54 22.8 0.87 0.23 14.8 0.60 0.41
+MoSca [38] 23.7 0.83 0.14 22.5 0.84 0.39 14.8 0.59 0.42
 SOM [70] 21.2 0.81 0.22 23.8 0.87 0.23 14.7 0.67 0.40
@@
 STG [42] 23.1 0.87 0.39 20.1 0.68 0.33 22.7 0.95 0.16
+MoSca [38] 20.9 0.86 0.61 16.7 0.54 0.57 22.1 0.94 0.21
 SOM [70] 18.7 0.84 0.41 17.1 0.65 0.41 23.8 0.95 0.18
@@
 STG [42] 24.5 0.97 0.10 20.2 0.77 0.37 11.1 0.51 0.44
+MoSca [38] 24.8 0.95 0.14 21.2 0.72 0.49 12.0 0.51 0.38
 SOM [70] 23.9 0.95 0.18 15.3 0.81 0.33 14.9 0.69 0.28
@@
 STG [42] 10.6 0.60 0.39 11.9 0.51 0.60 20.6 0.87 0.16
+MoSca [38] 22.2 0.87 0.12 19.9 0.81 0.12 17.1 0.78 0.32
 SOM [70] 13.9 0.68 0.34 17.4 0.79 0.14 20.7 0.85 0.17
@@
 STG [42] 19.6 0.86 0.19 16.8 0.66 0.43 24.7 0.90 0.17
+MoSca [38] 14.8 0.80 0.49 18.4 0.62 0.41 26.0 0.88 0.17
 SOM [70] 21.0 0.90 0.17 16.3 0.61 0.43 27.4 0.93 0.12
@@
 STG [42] 26.0 0.87 0.14 22.1 0.92 0.12 19.3 0.76 0.30
+MoSca [38] 15.3 0.56 0.50 18.3 0.86 0.25 19.5 0.76 0.34
 SOM [70] 16.3 0.60 0.24 22.5 0.93 0.15 19.3 0.80 0.26
-B.2 Performance breakdown: multiview setting
-We provide per-scene breakdown performance for the multiview video reconstruction setting in Table 10.
+E.4 Performance breakdown: multiview setting
+We provide per-scene breakdown performance for the multiview video reconstruction setting in Table G.
 We report results for D-3DGS [79], 4DGS [74], and STG [42], which serve as the most common and recent baselines for the DyNVS task.
-Note that since SOM [15] is specialized for monocular video setups, we omit its performance in the multiview evaluation.
-Table 10: Per-scene breakdown results for all 17 scenes under the multiview setting.
+Note that since MoSca [38] and SOM [70] are specialized for monocular video setups, we omit its performance in the multiview evaluation.
+Table G: Per-scene breakdown results for all 17 scenes under the multiview setting.
 Method Cereal Ship Hanok
@@
 STG [42] 16.9 0.62 0.36 23.8 0.93 0.11 21.0 0.79 0.30
-B.3 Additional Qualitative Results
-Figure 10 shows additional qualitative results, where all methods fail to accurately capture multi-body interactions, producing blurred or under-reconstructed outputs.
-For the 4DGS [37], using the default grid learning rate frequently causes NaN values in the loss.
-Reducing the learning rate stabilizes training but results in poor dynamic capture, with outputs resembling static scenes.
+E.5 Additional Qualitative Results
+Figure D and Figure E present qualitative results from monocular settings, revealing that all methods struggle to accurately capture multi-body interactions.
+This leads to common issues like needle-like artifacts and under-reconstruction of dynamic regions.
+For 4DGS [74], its default grid learning rate often causes NaN loss values; while reducing it stabilizes training, the output resembles static scenes due to poor dynamic capture.
 (a) GT image
@@
 (d) STG [42]
-(e) SOM [15]
-Figure 10: Qualitative results of recent DyNVS methods on the basin, bouncing balls, box-smoke, pancake, and ice scenes with monocular training setup.
+(e) MoSca [38]
+(f) SOM [70]
+Figure D: Qualitative results of recent DyNVS methods on the basin, bouncing balls, box-smoke, pancake, and ice scenes with monocular training setup.
 These results show that all methods frequently exhibit needle-like artifacts and fail to reconstruct dynamic elements accurately.
-Appendix C License
-Our dataset is released under the Creative Commons Attribution-NonCommercial (CC BY-NC) license.
+(a) GT image
+(b) D-3DGS [79]
+(c) 4DGS [74]
+(d) STG [42]
+(e) MoSca (2024)
+(f) SOM [70]
+Figure E: Qualitative results of recent DyNVS methods on the lucy, and hanok scenes with monocular setup.
+All methods struggle to accurately capture multi-body interactions by frequently exhibiting needle-like artifacts and failing to reconstruct dynamic elements accurately.
+F Limitations & Broad Impact
+Limitations
+While our PhysGaia is built upon a physics simulator to provide physically consistent supervision, it inherits the inherent approximations of the underlying physical models.
+In practice, physics simulators rely on simplified constitutive assumptions and numerical discretizations to make the problem tractable and stable.
+As a result, certain aspects of real-world dynamics, such as complex material responses or fine-scale interactions, may not be fully captured.
+We view this not as a limitation of our dataset alone, but as an inherent characteristic of simulation-based approaches, and an important direction for future work.
+Broad impact
+From a positive perspective, our work advances research on physically plausible 4D reconstruction.
+This, in turn, significantly improves monocular video reconstruction, an essential technology for future AR/VR applications.
+However, such progress in 4D reconstruction, especially given its potential for scene editing of existing videos, may raise intellectual property concerns regarding the original video content.
+G License
+Our benchmark is released under the Creative Commons Attribution-NonCommercial (CC BY-NC) license.
 The mesh used for the lucy scene is sourced from the Stanford 3D Repository, which permits usage for research purposes.
-Other mesh objects are licensed under CC-BY 4.0, as detailed in Table 11.
-For the texture maps in the textile categories, we utilized images from Pixabay, which are freely available with contributor consent and comparable to a CC-BY 4.0 license, also summarized in Table 11.
-Therefore, all data included in our dataset comply with usage rights and do not pose intellectual property issues.
-Table 11: License of the sources used for generating data.
+Other mesh objects are licensed under CC-BY 4.0, as detailed in Table H.
+For the texture maps in the textile categories, we utilized images from Pixabay, which are freely available with contributor consent and comparable to a CC-BY 4.0 license, also summarized in Table H.
+Therefore, all data included in our benchmark comply with usage rights and do not pose intellectual property issues.
+Table H: License of the sources used for generating data.
+The licenses for the mesh objects follow Creative Commons terms, except for the Lucy mesh sourced from the Stanford 3D Repository [37].
+The images used in the PhysGaia scenes, along with their download sources, are licensed under the Content License granted by Pixabay.
     Scene     Resource name     License     Access
@@
     Basin     Yourialka     Content License     Pixabay
-Appendix D Limitations & Broad Impact
-Limitations
-Although our PhysGaia dataset holds a crucial position in the DyNVS field by enabling accurate evaluation with ground-truth physical properties, it is inherently limited to synthetic scenarios.
-To bridge this gap towards more realistic applications, we believe domain adaptation approaches, such as those from Syn2Real [79], could be adopted.
-However, such methods must be carefully designed to maintain multi-view consistency across frames.
-We leave this important domain transfer for future work, confident that our dataset’s precise physical information will be invaluable for developing these advanced techniques.
-Broad impact
-From a positive perspective, our work advances research on physically plausible 4D reconstruction.
-This, in turn, significantly improves monocular video reconstruction, an essential technology for future AR/VR applications.
-However, such progress in 4D reconstruction, especially given its potential for scene editing of existing videos, may raise intellectual property concerns regarding the original video content.